### v4

# Variation of cones of divisors in a family of varieties – Fano type case

Abstract.
We investigate the relationship between the Fano type property on fibers over a Zariski dense subset and the global Fano type property. We establish the invariance of Néron-Severi spaces, nef cones, effective cones, movable cones, and Mori chamber decompositions for a family of Fano type varieties after a generically finite base change. Additionally, we show the uniform behavior of the minimal model program for this family. These results are applied to the boundedness problem of Fano type varieties.
Key words and phrases:
Fano type variety, Nef cone, Movable cone, Mori chamber decomposition.
2020 Mathematics Subject Classification:
14J45, 14E30
1. Introduction
We work over the field of complex numbers.
For a projective variety $X$, various cones of divisors can be associated within the Néron-Severi space $N^{1}(X)$. Notable among them are the effective cone ${\rm{Eff}}(X)$, the nef cone ${\rm{Nef}}(X)$, and the movable cone ${\rm{Mov}}(X)$. Moreover, the movable cone admits a finer decomposition known as the Mori chamber decomposition. These cones as well as their structures “linearize” the geometric properties of the variety.
Suppose that $X\to T$ is a family of varieties. It is natural to associate the aforementioned cones with each fiber and study how the deformation of the varieties affect the structure of the various cones. However, these cones may change drastically. For instance, there may not exist an open subset $U\subset T$ where the dimension of the Néron-Severi space $\dim N^{1}(X_{t})$ remains constant for all $t\in U$.
One motivation for this note is to apply the cone structure to study the moduli problem of Fano type varieties. Therefore, instead of considering a family of Fano type varieties, one should only assume the existence of a Zariski dense subset over which the fibers are of Fano type. In general, such a Zariski dense subset might be very sparse. Thus, it is natural to ask:
Question 1.1.
Let $X\to T$ be a family of varieties. Suppose that the fibers are of Fano type over a Zariski dense subset of $T$. After shrinking $T$, is $X$ of Fano type over $T$?
This natural question seems to be rather subtle, and we can only provide a partial answer.
Theorem 1.2.
Let $f:X\to T$ be a projective, surjective morphism between varieties such that $f_{*}\mathcal{O}_{X}=\mathcal{O}_{T}$. Suppose that there exists a Zariski dense subset $S\subset T$ such that fibers $X_{s},s\in S$ are Fano type varieties. Then, after shrinking $T$, there exists a divisor $D$ on $X$ such that $(X,D)$ has lc singularities and
$K_{X}+D\sim_{\mathbb{Q}}0/T.$
Furthermore, if $X_{s},s\in S$ fall into one of the following cases:
  1. (i)
    $X_{s}$ is a klt weak Fano variety for any $s\in S$,
  2. (ii)
    $X_{s}$ is of Fano type for any $s\in S$ and $S$ consists of very general points,
  3. (iii)
    $X_{s}$ is of Fano type for any $s\in S$ and $\{X_{s}\mid s\in S\}$ admits bounded klt complements,
then, after shrinking $T$, we have that $X$ is of Fano type over $T$.
By the invariance of plurigenera, varieties of general type deform to varieties of general type (see [HMX18]). The above theorem can be viewed as a weak analogy to this result for Fano type varieties. The proof of this natural theorem, however, requires deep results on complements as developed in [Birkar19]. With this result in mind, to study the deformation of cones of Fano type varieties in the above cases, we can always assume that each fiber is of Fano type after shrinking the base. It turns out that various cones exhibit nice behaviors in this family. As a first step, we show that, after a generically finite base change and a suitable shrinking of the base, the Néron-Severi space of each fiber coincides with the restriction of the Néron-Severi space of the total space.
Theorem 1.3.
Let $f:X\to T$ be a projective fibration. Suppose that $S\subset T$ is a Zariski dense subset such that for any $s\in S$, the fiber $X_{s}$ has rational singularities and
$H^{1}(X_{s},\mathcal{O}_{X_{s}})=H^{2}(X_{s},\mathcal{O}_{X_{s}})=0.$
  1. (1)
    If the natural restriction map $N^{1}(X/T)\to N^{1}(X_{t})$ is surjective for very general $t\in T$, then there exists a non-empty open subset $T_{0}\subset T$, such that ${\mathcal{P}}ic_{X_{T_{0}}/T_{0}}\otimes\mathbb{R}$ is a constant sheaf in the Zariski topology.
  2. (2)
    Up to a generically finite base change of $T$, there exists a non-empty open subset $T_{0}\subset T$, such that ${\mathcal{P}}ic_{X_{T_{0}}/T_{0}}\otimes\mathbb{R}$ is a constant sheaf in the Zariski topology.
Moreover, in both of the above cases, for any open subset $U\subset T_{0}$, the natural restriction maps
$N^{1}(X_{T_{0}}/T_{0})\to N^{1}(X_{U}/U)\to N^{1}(X_{t})$
are isomorphisms for any $t\in U$.
In the above theorem, “up to a generically finite base change of TT” means that we take a generically finite morphism $T^{\prime}\to T$ and replace $T$ with $T^{\prime}$. The non-empty open subset $T_{0}$ is then taken inside this new base $T^{\prime}$. Hence, after possibly further shrinking $T_{0}$, we are indeed performing an étale base change of an open subset of the original base.
The proof of the above result follows from an argument akin to [dFH11, §6], where a similar result is established in a slightly different setting. Due to the monodromy phenomenon, even in a locally trivial family, there are no natural identifications between the Néron-Severi spaces of different fibers. Theorem 1.3 enables comparisons between these spaces and their associated cones.
The following fiber-wise small $\mathbb{Q}$-factorial modification illustrates the necessity of considering families whose fibers have rational singularities and satisfy $h^{i}(X_{s},\mathcal{O}_{X_{s}})=0$ for $i=1,2$, rather than restricting only to varieties of Fano type. Indeed, these properties are preserved under resolutions, which may result in varieties that are no longer of Fano type. Such fiber-wise small $\mathbb{Q}$-factorial modifications are essential in establishing the boundedness of moduli problems, although they have been used in the literature without proper justification.
Theorem 1.4.
Let $f:X\to T$ be a projective fibration. Suppose that $S\subset T$ is a Zariski dense subset such that for any $s\in S$, the fiber $X_{s}$ has rational singularities and
$H^{1}(X_{s},\mathcal{O}_{X_{s}})=H^{2}(X_{s},\mathcal{O}_{X_{s}})=0.$
Assume that there exists a divisor $\Delta$ such that $(X,\Delta)$ has klt singularities. Then, up to a generically finite base change of $T$, there exist a birational morphism $Y\to X$ and a non-empty open subset $T_{0}\subset T$ such that $Y\to X$ and each fiber $Y_{t}\to X_{t}$ for $t\in T_{0}$ are small $\mathbb{Q}$-factorial modifications.
Moreover, for any open subset $U\subset T_{0}$, the natural restriction maps
$\begin{split}&N^{1}(Y_{T_{0}}/T_{0})\to N^{1}(Y_{U}/U)\to N^{1}(Y_{t})\\ &N^{1}(X_{T_{0}}/T_{0})\to N^{1}(X_{U}/U)\to N^{1}(X_{t})\end{split}$
are isomorphisms for any $t\in U$.
The following result establishes the generic deformation invariance of various cones up to a base change.
Theorem 1.5.
Let $f:X\to T$ be a projective fibration. Suppose that there exists a Zariski dense subset $S\subset T$ such that the fibers $X_{s},s\in S$ satisfy one of the cases in Theorem 1.2.
  1. (1)
    If the natural restriction map $N^{1}(X/T)\to N^{1}(X_{t})$ is surjective for very general $t\in T$, then there exists a non-empty open subset $T_{0}\subset T$, such that the Néron-Severi space $N^{1}(X_{t})$ as well as the cones ${\rm{Eff}}(X_{t})$, ${\rm{Nef}}(X_{t})$ and ${\rm{Mov}}(X_{t})$ are deformation invariant on $T_{0}$. Moreover, the Mori chamber decomposition on ${\rm{Mov}}(X_{t})$ is also deformation invariant on $T_{0}$.
  2. (2)
    Up to a generically finite base change of $T$, there exists a non-empty open subset $T_{0}\subset T$, such that the Néron-Severi space $N^{1}(X_{t})$ as well as the cones ${\rm{Eff}}(X_{t})$, ${\rm{Nef}}(X_{t})$ and ${\rm{Mov}}(X_{t})$ are deformation invariant on $T_{0}$. Moreover, the Mori chamber decomposition on ${\rm{Mov}}(X_{t})$ is also deformation invariant on $T_{0}$.
For the precise definition of the Mori chamber decomposition, see (5.3). Theorem 1.5 is a combination of Proposition 5.6, Theorem 4.5 (also see Corollary 4.7), and Proposition 5.7, where the claims are proved under some more general setting.
Shrinking bases is a generally allowed operation in the moduli problem, as we can apply Noetherian induction. Moreover, this operation is necessary even if $X\to T$ is a family of $\mathbb{Q}$-factorial terminal Fano varieties (see [Totaro12]). On the other hand, the nef cones are indeed locally constant (i.e., deformation invariant) when $f:X\to T$ is a smooth family of Fano manifolds (see [Wis91, Wis09]). When $X\to T$ is a family of $\mathbb{Q}$-factorial terminal Fano varieties over a curve, [dFH11] showed that the movable cones are locally constant. Under the same setting, the local constancy of the Mori chamber decompositions is also established for special families of Fano varieties, notably when $\dim(X/T)\leq 3$ (see counterexamples in [Totaro12] for cases where $\dim(X/T)>3$). Moreover, under various restrictions, the invariance of effective cones and movable cones is established in [HX15]. [Sho20, §4] established similar results when $X$ is of Fano type over $T$. In fact, [Sho20, §4] showed that the sets of integral divisors associated with various cones remain invariant. The case of effective cones is also addressed in a recent preprint [CHHX25]. These results are indispensable in the theory of complements.
Next, we explore the uniform behavior of the minimal model program (MMP) for a family of varieties. In the case where the fibers are surfaces, the stability of (1)(-1)-curves and the simultaneous blowing down of (1)(-1)-curves have long been established (see [Kod63, Iit70]). We demonstrate the good behavior of the minimal model program for families of Fano type varieties. As a direct consequence, we obtain the deformation invariance of the Mori chamber decompositions of the movable cones. This circle of ideas has already been applied in [dFH11, HX15]. In the general setting, we establish the following deformation invariance of the MMP.
Theorem 1.6.
Let $f:X\to T$ be a projective fibration. Suppose that there exists a Zariski dense subset $S\subset T$ such that the fibers $X_{s}$ for $s\in S$ satisfy one of the cases in Theorem 1.2. Then, after a generically finite base change of $T$, there exists a non-empty Zariski open subset $T_{0}\subset T$ such that any Zariski open subset $U\subset T_{0}$ satisfies the following properties.
  1. (1)
    Suppose that $X^{\prime}/U\in{\rm{bc}}(X_{U}/U)$. If $g:X^{\prime}\dashrightarrow Y/U$ is a birational contraction with $Y$ a $\mathbb{Q}$-factorial variety, then for any $t\in U$, the induced map $g_{t}:X^{\prime}_{t}\dashrightarrow Y_{t}$ is again a birational contraction. Moreover, for any $\mathbb{R}$-Cartier divisor $\mathcal{D}$ on $X^{\prime}$, we have
    $(g_{t})_{*}(\mathcal{D}|_{X^{\prime}_{t}})\sim_{\mathbb{R}}(g_{*}\mathcal{D})|_{Y_{t}},\quad t\in U.$
  2. (2)
    For any $\mathbb{R}$-divisor $\mathcal{D}\in N^{1}(X_{U}/U)$, a sequence of $\mathcal{D}$-MMP/U/U induces a sequence of $\mathcal{D}|_{X_{t}}$-MMP of the same type for each $t\in U$.
  3. (3)
    Conversely, any $\mathbb{R}$-divisor $D\in N^{1}(X_{t})$ with $t\in U$, a sequence of $D$-MMP on $X_{t}$ is induced by a sequence of $\mathcal{D}$-MMP/U/U on $X_{U}/U$ of the same type. Moreover, we can choose $\mathcal{D}$ such that $\mathcal{D}|_{X_{t}}\sim_{\mathbb{R}}D$.
  4. (4)
    If $Y/U\in{\rm{bc}}(X_{U}/U)$, then the natural maps
    $N^{1}(Y/U)\to N^{1}(Y_{t}),\quad{\rm{Nef}}(Y/U)\to{\rm{Nef}}(Y_{t}),\quad{\rm{NE}}(Y_{t})\to{\rm{NE}}(Y/U)$
    are isomorphisms for any $t\in U$.
In the above theorem, ${\rm{bc}}(X_{U}/U)$ consists of all birational contractions from $X_{U}/U$. See Section 5.1 for the formal definition.
As an immediate application of these results, we establish a boundedness result concerning Fano type varieties. Let $X\to T$ be a fixed family of Fano type varieties. Let
$\mathcal{S}\coloneqq\{X_{t}\mid t\in T\}$
be a set consisting of closed fibers. For any $X_{t}\in\mathcal{S}$, let ${\rm bcm}(X_{t})$ consist of normal projective varieties which are birational contraction models of $X_{t}$, that is,
${\rm bcm}(X_{t})\coloneqq\{Y\mid\text{there exists a birational contraction~}X_{t}\dashrightarrow Y\}.$
Set ${\rm bcm}(\mathcal{S})\coloneqq\cup_{X_{t}\in\mathcal{S}}{\rm bcm}(X_{t})$. Recall that a set of varieties is bounded if it can be parametrized by the fibers of a morphism between two schemes of finite type. We have the following result on boundedness of ${\rm bcm}(\mathcal{S})$:
Theorem 1.7.
The set ${\rm bcm}(\mathcal{S})$ is bounded.
Theorem 1.7 is proved following the guiding principle that boundedness of varieties can often be deduced from birational boundedness together with the finiteness of models. This strategy has been successfully applied to many boundedness problems (see [HX15, HMX18, MST20, FHS20], etc.). The proof of Theorem 1.7 is essentially a restatement of the argument of [HX15, Theorem 1.3], with additional clarification on how to obtain a fiber-wise $\mathbb{Q}$-factorial modification.
Theorems 1.5, 1.6, and 1.7 are further generalized and applied in the subsequent paper [CLLZ25], where a hybrid of Mori dream spaces and Calabi-Yau type varieties (termed Morrison-Kawamata dream spaces) is investigated.
We outline the structure of the paper. Section 2 provides the necessary background and sets up notation. Section 3 is devoted to the proof of Theorem 1.2, which gives a partial answer to Question 1.1. Section 4 investigates the deformation invariance of Néron-Severi spaces and nef cones. Section 5 establishes fiber-wise small $\mathbb{Q}$-factorial modifications, the uniform behavior of the MMP for families of Fano type varieties, and the deformation invariance of effective cones and Mori chamber decompositions. Finally, Section 6 presents an application of the preceding results to the boundedness of birational contraction models of Fano type varieties.
Acknowledgements
We would like to thank Christopher Hacon, Xiaowen Hu, Yifei Zhu and Ziquan Zhuang for answering our questions, and Guodu Chen for helpful discussions. We are also grateful to Stefano Filipazzi for helping us clarify the importance of taking base changes, and to Vyacheslav Shokurov for bringing our attention to his paper [Sho20]. S. Choi is partially supported by Samsung Science and Technology Foundation under Project Number SSTF-BA2302-03. Z. Li is partially supported by NSFC (No.12471041), the Guangdong Basic and Applied Basic Research Foundation (No.2024A1515012341), and a grant from SUSTech. C. Zhou is supported by a grant from Xiamen University (No. X2450214).
2. Preliminaries
We introduce the necessary background materials. Along with this process, we fix the notation and terminologies.
A variety means an integral separated scheme of finite type over $\mathbb{C}$. A point of a variety is understood to be a closed point and an open subset of a variety is meant to be a non-empty Zariski open subset, unless explicitly stated otherwise. For a variety $T$, by ”shrinking TT” we mean replacing $T$ with a Zariski open subset. A subset $S$ of a variety consists of very general points if there exist at most countably many non-empty open subsets $U_{i}$ such that $S=\cap U_{i}$.
A projective morphism $f:X\to S$ between normal varieties is called a fibration if it is surjective and $f_{*}\mathcal{O}_{X}=\mathcal{O}_{S}$. If $S^{\prime}\to S$ is a morphism, then $X_{S^{\prime}}$ denotes the fiber product $X\times_{S}S^{\prime}$. Similarly, if $\mathcal{D}$ is an $\mathbb{R}$-Cartier divisor on $X$, then $\mathcal{D}_{S^{\prime}}$ denotes the pullback of $\mathcal{D}$ to $X_{S^{\prime}}$. A birational map $g:X\dasharrow X^{\prime}$ between normal varieties is called a birational contraction if $g^{-1}$ does not contract any divisor.
Suppose that $\Delta\geq 0$ is an $\mathbb{R}$-divisor on a normal variety $X$. Then $(X,\Delta)$ is called a log pair. A log pair $(X,\Delta)$ has klt singularities if $K_{X}+\Delta$ is $\mathbb{R}$-Cartier where $K_{X}$ is the canonical divisor of $X$, and there exists a log resolution $\pi:Y\to X$ such that in the expression
$K_{Y}=\pi^{*}(K_{X}+\Delta)+D,$ (2.1)
the coefficients of $D$ are greater than 1-1. Note that in (2.1), $K_{Y}$ is chosen to be the unique Weil divisor on $Y$ such that $\pi_{*}K_{Y}=K_{X}$. Similarly, if the coefficients of $D$ are greater than or equal to 1-1, then $(X,\Delta)$ is said to have lc singularities. See [KM98, §2.3] for more detailed discussions.
Let $f:X\to T$ be a projective morphism between normal varieties. We use “/T/T” to denote properties which are relative to $T$. Then $X$ is of Fano type/T/T if there exists a divisor $\Delta$ such that $(X,\Delta)$ has klt singularities and $(K_{X}+\Delta)$ is ample/T/T. Note that in the definition, we do not assume that $K_{X}$ is a $\mathbb{Q}$-Cartier divisor. By passing to a small $\mathbb{Q}$-factorization, it is straightforward to see that in the definition of Fano type variety, $\Delta$ can be chosen to be a $\mathbb{Q}$-divisor. Moreover, if $K_{X}$ is a $\mathbb{Q}$-Cartier divisor, then $X$ is of Fano type/T/T if and only if there exists a $\mathbb{Q}$-Cartier divisor $B$ which is big over $T$ such that $(X,B)$ has klt singularities and $K_{X}+B\sim_{\mathbb{Q}}0/T$. Besides, $X\to T$ is called a family of Fano type varieties if all of its closed fibers are Fano type varieties. A variety $X$ is called weak Fano if $K_{X}$ is nef and big. Hence, a klt weak Fano variety is automatically of Fano type. For some $n\in\mathbb{N}$, a divisor $\Delta\in|-nK_{X}|$ is called an lc (resp. klt) $n$-complement of $K_{X}$ if $(X,\frac{1}{n}\Delta)$ has lc (resp. klt) singularities (see [PS09] for a more general definition). A set of varieties $\mathcal{P}$ is said to admit bounded lc (resp. klt) complements if there exists an integer $n$, depending only on $\mathcal{P}$, such that every element of $\mathcal{P}$ has an lc (resp. klt) $n$-complement.
A Cartier divisor $D$ is movable/T/T on $X/T$ if the codimension of the support of the sheaf ${\rm coker}(f^{*}f_{*}\mathcal{O}_{X}(D)\to\mathcal{O}_{X}(D))$ is greater than or equal to $2$. We list relevant vector spaces and cones which appear in this paper:
  1. (1)
    ${\rm{Pic}}(X/T)_{\mathbb{Z}}$: the abelian group generated by Cartier divisors on $X$ modulo linear equivalences over $T$.
  2. (2)
    ${\rm{Pic}}(X/T)$: the $\mathbb{R}$-vector space generated by Cartier divisors on $X$ modulo linear equivalences over $T$, which is the same as ${\rm{Pic}}(X/T)_{\mathbb{Z}}\otimes_{\mathbb{Z}}\mathbb{R}$ .
  3. (3)
    $N^{1}(X/T)$: the $\mathbb{R}$-vector space generated by Cartier divisors on $X$ modulo numerical equivalences over $T$.
  4. (4)
    $N_{1}(X/T)$: the $\mathbb{R}$-vector space generated by curves that are contracted by $X\to T$ modulo numerical equivalences.
  5. (5)
    ${\rm{NE}}(X/T)\subset N_{1}(X/T)$: the cone generated by curves that are contracted by $X\to T$ (its closure $\overline{{\rm{NE}}}(X/Y)$ is called the Mori cone).
  6. (6)
    (nef cone) ${\rm{Nef}}(X/T)\subset N^{1}(X/T)$: the cone generated by nef Cartier divisors on $X$ over $T$.
  7. (7)
    (effective cone) ${\rm{Eff}}(X/T)\subset N^{1}(X/T)$: the cone generated by effective Cartier divisors on $X$ over $T$ (its closure $\overline{\rm{Eff}}(X/T)$ is called the pseudo-effective cone).
  8. (8)
    (movable cone) ${\rm{Mov}}(X/T)\subset N^{1}(X/T)$: the cone generated by movable Cartier divisors on $X$ over $T$.
If $T={\rm{Spec}}(\mathbb{C})$, then we will omit $T$ in the above notations. For simplicity, when we say that $D\in N^{1}(X/T)$ is a divisor, we mean that $D]\in N^{1}(X/T)$ with $D$ an $\mathbb{R}$-Cartier divisor. This convention applies to other cones as well. We use ${\rm Int}(C)$ to denote the relative interior of the cone $C$. A cone $C\subset N^{1}(X/T)$ is called a rational polyhedral cone if it is generated by finitely many rational rays (hence, it is a closed cone). In general, ${\rm{Pic}}(X/T)$ may be an infinite dimensional vector space. However, when $X/T$ is of Fano type, it is a finite-dimensional vector space. Besides, many of the above cones are neither open nor closed. Even worse, their closures may not be generated by rational rays. In the literature, there are various conventions regarding effective and movable cones (e.g., taking the closure of the cone or selecting rational elements within the cones). However, when $X/T$ is of Fano type, ${\rm{NE}}(X/T),{\rm{Nef}}(X/T),{\rm{Eff}}(X/T)$ and ${\rm{Mov}}(X/T)$ are all closed rational polyhedral cones. Since this paper focuses on Fano type varieties, we do not need to distinguish these variations on the definitions of cones.
3. Families with fibers of Fano type varieties over Zariski dense subsets
The purpose of this section is to prove Theorem 1.2. We begin by establishing a preparatory lemma.
Lemma 3.1.
Let $X\to T$ be a projective fibration. Assume that
  1. (1)
    there exists a Zariski dense subset $S\subset T$ such that $X_{s}$ is of Fano type for each $s\in S$, and
  2. (2)
    there exists $\Delta\geq 0$ such that $(X,\Delta)$ is lc and $K_{X}+\Delta\sim_{\mathbb{Q}}0/T$.
Then, after shrinking $T$, there exists a small $\mathbb{Q}$-factorial modification $Z\to X$ such that $Z$ has klt singularities.
Proof.
Let $h:W\to X$ be a $\mathbb{Q}$-factorial dlt modification of $(X,\Delta)$ (see [BCHM10]) such that
$K_{W}+\Delta_{W}+{\rm{Exc}}(h)=h^{*}(K_{X}+\Delta),$
where $\Delta_{W}$ is the strict transform of $\Delta$, and ${\rm{Exc}}(h)$ is the union of $h$-exceptional divisors (with coefficient $1$). Hence, any $h$-exceptional divisor is an lc place of $(X,\Delta)$. Run a $(K_{W}+{\rm{Exc}}(h))$-MMP/$X$,
$\phi:W\dasharrow Z/X,$
which terminates at $Z$ with $K_{Z}+\phi_{*}{\rm{Exc}}(h)$ nef over $X$ (see [HX13]). We claim that $Z\to X$ is a small morphism after shrinking $T$.
This is straightforward if $K_{X}$ is $\mathbb{Q}$-Cartier: as $X_{s}$ is of Fano type for $s\in S$, we see that $X_{s}$ has klt singularities. Therefore, $X$ still has klt singularities after shrinking $T$. This means that the lc centers of $(X,\Delta)$ are contained in ${\rm{Supp}}(\Delta)$. Hence, we have $\Delta_{W}+E=h^{*}\Delta$ with ${\rm{Supp}}(E)={\rm{Exc}}(h)$. This implies that
$K_{W}+{\rm{Exc}}(h)=h^{*}K_{X}+E,$
and any $(K_{W}+{\rm{Exc}}(h))$-MMP/$X$ will contract $E$. In particular, $Z\to X$ is a small morphism. In the sequel, we explain that the same property still holds even if $K_{X}$ is not $\mathbb{Q}$-Cartier.
[…]
Suppose that we have $\phi_{*}{\rm{Exc}}(h)\neq 0$. As $h({\rm{Exc}}(h))\subset{\rm{Supp}}(\Delta)$, we have $q(\phi_{*}{\rm{Exc}}(h))\subset{\rm{Supp}}(\Delta)$, where $q:Z\to X$. We can cut $X$ by sufficiently ample and general hypersurfaces $H_{i}\subset X,i=1,\ldots,l$ until
$q\left(\phi_{*}{\rm{Exc}}(h)\right)\cap H_{1}\cap\cdots\cap H_{l}$
consists of just points. We use the same notation to denote the varieties and the divisors after this cutting. Note that as these hypersurfaces are sufficiently ample, $K_{Z}+\phi_{*}{\rm{Exc}}(h)$ is still nef over $X$. Choose any $x_{0}\in q\left(\phi_{*}{\rm{Exc}}(h)\right)\cap H_{1}\cap\cdots\cap H_{l}$. Then $q^{-1}(x_{0})$ is of positive dimension. To be precise, we write
$q^{-1}(x_{0})=F\cup F^{\prime},$
where $F$ consists of $q$-exceptional divisors, and $F^{\prime}$ consists of higher codimensional subvarieties. As $x_{0}\in q(\phi_{*}{\rm{Exc}}(h))\subset{\rm{Supp}}(\Delta)$, and $q$ has connected fibers, we see that
$q^{-1}(x_{0})\cap{\rm{Supp}}(\Delta_{Z})\neq\emptyset.$
Thus, there exists a curve $\ell\subset q^{-1}(x_{0})$ such that
${\rm{Supp}}(\Delta_{Z})\cap\ell\neq\emptyset\quad\text{and}\quad\ell\not\subset{\rm{Supp}}(\Delta_{Z}).$
This is obvious if $F^{\prime}\not\subset{\rm{Supp}}(\Delta_{Z})$. If $F^{\prime}\subset{\rm{Supp}}(\Delta_{Z})$, then as ${\rm{Supp}}(\Delta_{Z})$ does not contain $q$-exceptional divisors and $q^{-1}(x_{0})$ is connected, we can find a curve $\ell\subset F$ satisfies the desired property. However, as
$K_{Z}+\Delta_{Z}+\phi_{*}{\rm{Exc}}(h)=q^{*}(K_{X}+\Delta),$
and $K_{Z}+\phi_{*}{\rm{Exc}}(h)$ is nef/X/X, we have
$0<(K_{Z}++\phi_{*}{\rm{Exc}}(h))\cdot\ell+\Delta_{Z}\cdot\ell=q^{*}(K_{X}+\Delta)\cdot\ell=0.$
This is a contradiction. Therefore, $Z\to X$ is a small morphism. This shows that $Z\to X$ is a small $\mathbb{Q}$-factorial modification of $X$ and $Z$ has klt singularities. ∎
We can now prove Theorem 1.2 which is a partial answer to Question 1.1.
Proof of Theorem 1.2.
We proceed with the argument in several steps. In Steps 1 and 2, we apply uniform modifications to all three cases and then treat each case separately. In what follows, after shrinking $T$, we still use $S$ to denote the Zariski dense subset $S\cap T$.
Step 1. As each $X_{s},s\in S$ is a normal variety, it is regular in codimension $1$ and satisfies Serre’s condition $S_{2}$. Shrinking $T$, we can assume that $X$ is regular in codimension $1$. By [EGAIV, (9.9.2)(viii)], the set of points (not necessarily closed points)
$E_{2}\coloneqq\{p\mid\mathcal{O}_{X}|_{X_{p}}\text{~is~}S_{2}\}$
is a constructible set. As $S$ is Zariski dense and $S\subset E_{2}$, the generic point of $T$ lies in $E_{2}$. Therefore, after shrinking $T$, we can assume that $X$ is $S_{2}$ and $T$ is normal. This shows that $X$ is normal, and thus $f$ is a projective contraction.
Step 2. In this step, we show that in each of the three cases, $X$ can be further assumed to be $\mathbb{Q}$-factorial with klt singularities. More precisely, assuming that there exists a Zariski dense subset $S\subset T$ such that $X_{s}$ is of Fano type for each $s\in S$ (which is automatically satisfied in the three cases), we prove that $X$ admits a small $\mathbb{Q}$-factorial modification with klt singularities after shrinking $T$. Note that in this step, we only need to assume that $S$ is Zariski dense.
First, we claim that there exist a universal $m\in\mathbb{Z}_{>0}$ and a Weil divisor $B^{s}\in|-mK_{X_{s}}|$ such that $(X_{s},\frac{1}{m}B^{s})$ has lc singularities for any $s\in S$. As $X_{s}$ is of Fano type, we can replace it with a small $\mathbb{Q}$-factorial modification $Y_{s}\to X_{s}$. Then $Y_{s}$ is still of Fano type, and it suffices to show the claim for $Y_{s}$. Therefore, we can assume that $X_{s}$ is $\mathbb{Q}$-factorial. Runing a $(-K_{X_{s}})$-MMP, $X_{s}\dasharrow X^{\prime}_{s}$, we have that $X^{\prime}_{s}$ is a Fano type variety with $K_{X^{\prime}_{s}}$ nef. By [Birkar19, Theorem 1.7], there exist a universal $m\in\mathbb{Z}_{>0}$ and a Weil divisor $D^{s}\in|-mK_{X^{\prime}_{s}}|$ such that $(X^{\prime}_{s},\frac{1}{m}D^{s})$ has lc singularities. Let $p:W\to X_{s},q:W\to X^{\prime}_{s}$ be a resolution of $X_{s}\dasharrow X^{\prime}_{s}$. Then we have
$mp^{*}K_{X_{s}}=mq^{*}K_{X_{s}^{\prime}}-F$
for some $q$-exceptional divisor $F\geq 0$. Hence, we have
$mp^{*}K_{X_{s}}+F+q^{*}D^{s}=q^{*}(mK_{X_{s}^{\prime}}+D^{s})\sim 0.$
In particular, $F+q^{*}D^{s}$ is an effective Weil divisor. Set
$B^{s}=p_{*}(F+q^{*}D^{s}).$
We have $B^{s}\in|-mK_{X_{s}}|$. As
$p^{*}(mK_{X_{s}}+B^{s})=q^{*}(mK_{X_{s}^{\prime}}+D^{s})\sim 0,$
we see that $(X_{s},\frac{1}{m}B^{s})$ still has lc singularities.
Next, we show that there exists a divisor $B$ such that $(X,\frac{1}{m}B)$ is lc and $mK_{X}+B\sim 0/T$. Consider the coherent sheaf $\mathcal{O}_{X}(-mK_{X})$. By generic flatness, we can assume that $\mathcal{O}_{X}(-mK_{X})$ is flat over $T$ after shrinking $T$. Replacing $T$ by an open affine subset, we can assume that
$H^{0}(T,f_{*}\mathcal{O}_{X}(-mK_{X}))\to H^{0}(X_{t},\mathcal{O}_{X}(-mK_{X})|_{X_{t}})$
is surjective for any $t\in T$. This uses [Har77, Chapter III, Theorem 12.8, Corollary 12.9], and the fact that the coherent sheaf $f_{*}\mathcal{O}_{X}(-mK_{X})$ is globally generated over an affine set. Shrinking $T$ further, we have $\mathcal{O}_{X}(-mK_{X})|_{X_{t}}=\mathcal{O}_{X_{t}}(-mK_{X_{t}})$ for each $t\in T$. In summary, we obtain a surjection
$H^{0}(T,f_{*}\mathcal{O}_{X}(-mK_{X}))\to H^{0}(X_{t},\mathcal{O}_{X_{t}}(-mK_{X_{t}})),\quad t\in T.$ (3.1)
As there exists $B^{s}\in|-mK_{X_{s}}|$ such that $(X_{s},\frac{1}{m}B^{s})$ is lc for any $s\in S$. By (3.1), let $B\in|-mK_{X}|$ be the element which maps to $B^{s}$, then $(X,\frac{1}{m}B)$ is lc in a Zariski neighborhood of $X_{s}$ (see [Laz04b, Theorem 9.5.19]). Shrinking $T$, we have that $(X,\frac{1}{m}B)$ is lc and $mK_{X}+B\sim 0/T$. This shows the first part of the claim.
To complete the remaining part of the claim, we explain that it suffices to assume that $X$ is a $\mathbb{Q}$-factorial klt variety. By Lemma 3.1, after shrinking $T$, $X$ admits a small $\mathbb{Q}$-factorial modification $W\to X$ such that $W$ has klt singularities. Shrinking $T$ further, we may assume that $W_{t}\to X_{t}$ is a small modification for each $t\in T$. In particular, the three cases remain valid after replacing $X$ with $W$. Moreover, if $W$ is of Fano type over $T$, then $X$ is of Fano type over $T$. Replacing $X$ by $W$, we can assume that $X$ is a $\mathbb{Q}$-factorial klt variety.
Step 3. In this step, we show the claim in Case (ii). Just as (3.1), for each $m\in\mathbb{N}$, there exists a non-empty open affine subset $U_{m}\subset T$ such that
$H^{0}(U_{m},f_{*}\mathcal{O}_{X}(-mK_{X}))\to H^{0}(X_{t},\mathcal{O}_{X}(-mK_{X_{t}}))$
is surjective for any $t\in U_{m}$. As $S$ consists of very general points, we have $S\cap(\cap_{m\in\mathbb{N}}U_{m})\neq\emptyset$. Therefore, if $s\in S\cap(\cap_{m\in\mathbb{N}}U_{m})$, then for any $m\in\mathbb{N}$, there exists an open subset $U_{m}$ such that the natural map
$H^{0}(U_{m},f_{*}\mathcal{O}_{X}(-mK_{X}))\to H^{0}(X_{t},\mathcal{O}_{X}(-mK_{X_{s}}))$
is surjective. As $X_{s}$ is of Fano type, there exists some $m\in\mathbb{N}$ and a divisor $D\in|-mK_{X_{s}}|$ such that $(X_{s},\frac{1}{m}D)$ has klt singularities. By the above surjection, there exists a divisor $\mathcal{D}\in H^{0}(X_{U_{m}},\mathcal{O}_{X}(-mK_{X}))$ such that $\mathcal{D}|_{X_{s}}=D$. As $(X_{s},\frac{1}{m}D)$ is klt, $(X,\frac{1}{m}\mathcal{D})$ is klt in a Zariski neighborhood of $X_{s}$ (see [Laz04b, Theorem 9.5.19]). Shrinking $T$, we can assume that $(X,\frac{1}{m}\mathcal{D})$ is klt. As $X$ is $\mathbb{Q}$-factorial and $X_{s}$ is of Fano type, we have that $\mathcal{D}_{s}$ is big for any $s\in S$. As $S$ consists of very general points, $\mathcal{D}$ is big over $T$ (see [Li22, Theorem 3.18]). Therefore, $X$ is of Fano type over $T$.
Step 4. In this step, we show that Case (i) can be reduced to Case (iii). As $X$ is $\mathbb{Q}$-factorial, shrinking $T$, there exists $n\in\mathbb{N}$ such that $nK_{X_{s}}$ is Cartier for each $s\in S$. As $K_{X_{s}}$ is nef and big, by the effective base-point free theorem ([Kol93]), there exists $m\in\mathbb{N}$ such that $|-mK_{X_{s}}|$ is base-point free for each $s\in S$. Since the claim holds when $\dim X_{s}=1$ (i.e., $X_{s}\simeq\mathbb{P}^{1}$), we can assume that $\dim X_{s}\geq 2$. Moreover, without loss of generality, we can also assume that $m\geq 2$. As $K_{X_{s}}$ is big and $\dim X_{s}\geq 2$, the base-point free linear system $|-mK_{X_{s}}|$ is not composed of a pencil. By Bertini theorem, a general element $B\in|-mK_{X_{s}}|$ is such that the pair $(X,\frac{1}{m}B)$ has klt singularities. This shows that $\{X_{s}\mid s\in S\}$ admits bounded klt complements.
Step 5. In this step, we establish the claim for Case (iii), which consequently implies the result for Case (i) as well.
Recall that $X$ is a $\mathbb{Q}$-factorial variety with klt singularities by Step 2. By assumption, there exists an integer $n\in\mathbb{N}$, such that each $X_{s},s\in S$ admits a klt $n$-complement. Same as (3.1), possibly shrinking $T$, the natural restriction map
$H^{0}(X,\mathcal{O}_{X}(-nK_{X}))\to H^{0}(X_{s},\mathcal{O}_{X_{s}}(-nK_{X_{s}}))$
is surjective for each $s\in S$. Therefore, if $\Delta_{s}\in|-nK_{X_{s}}|$ such that $(X_{s},\frac{1}{n}\Delta_{s})$ has klt singularities, then there exists $\Delta\in|-nK_{X}|$ such that $(X,\frac{1}{n}\Delta)$ has klt singularities in a Zariski neighborhood of $X_{s}$ (see [Laz04b, Theorem 9.5.19]). Shrinking $T$, we can assume that $(X,\frac{1}{n}\Delta)$ has klt singularities with $K_{X}+\frac{1}{n}\Delta\sim_{\mathbb{Q}}0$. Hence, there exists $\epsilon\in\mathbb{Q}_{>0}$ such that $(X,(\frac{1}{n}+\epsilon)\Delta)$ still has klt singularities. By [HMX18, Theorem 4.2], after shrinking $T$, we know that the Kodaira dimension $\kappa(X_{t},K_{X_{t}}+(\frac{1}{n}+\epsilon)\Delta_{t})$ is independent of $t\in T$. As $K_{X}+(\frac{1}{n}+\epsilon)\Delta\sim_{\mathbb{Q}}\epsilon n(-K_{X})$, we see that $K_{X}$ is big over $T$. This implies that $X$ is of Fano type over $T$. ∎
In analogy to Question 1.1, we ask:
Question 3.2.
Let $f:X\to T$ be a projective fibration. Suppose that there exists a Zariski dense subset $S\subset T$ such that $X_{s}$ is a Mori dream space for each $s\in S$. Is it true that $X$ is a Mori dream space over $T$ after shrinking $T$?
4. Deformation of Néron-Severi spaces and nef cones
In this section, we study the deformation of Néron-Severi spaces and various cones for a family of varieties under certain conditions.
4.1. Deformation of Néron-Severi spaces
Let $f:X\to T$ be a projective morphism such that $f_{*}\mathcal{O}_{X}=\mathcal{O}_{T}$. Let ${\mathcal{P}}ic_{X/T}$ be the sheaf associated to the relative Picard functor
$S\mapsto{\rm{Pic}}(X_{S})_{\mathbb{Z}}/{\rm{Pic}}(S)_{\mathbb{Z}}={\rm{Pic}}(X_{S}/S)_{\mathbb{Z}},$
where $S\subset T$ is a Zariski open subset. See [Kle05, §9.2] for details. Note that ${\mathcal{P}}ic_{X/T}$ is denoted by ${\rm{Pic}}_{(X/T)(\rm{zar})}$ in [Kle05, Definition 9.2.2]. In general, ${\mathcal{P}}ic_{X/T}(U)$ may not be ${\rm{Pic}}(X_{U}/U)_{\mathbb{Z}}$ for an open subset $U\subset T$ because of the sheafification. In fact, by [Kle05, (9.2.11.2)], we always have
${\mathcal{P}}ic_{X/T}(U)=H^{0}(U,R^{1}f_{*}\mathcal{O}^{*}_{X_{U}}).$ (4.1)
If $X\to T$ is a flat morphism, then by [Kle05, (9.2.11.3)], we have the exact sequence
$0\to{\rm{Pic}}(U)_{\mathbb{Z}}\to{\rm{Pic}}(X_{U})_{\mathbb{Z}}\to{\mathcal{P}}ic_{X/T}(U).$
In particular, we have the inclusion ${\rm{Pic}}(X_{U}/U)_{\mathbb{Z}}\hookrightarrow{\mathcal{P}}ic_{X/T}(U)$.
In the following, we show that if $X\to T$ is a family of Fano type varieties, then there exists a generically finite base change $T^{\prime}\to T$ and an open subset $U\subset T^{\prime}$, such that ${\mathcal{P}}ic_{X_{U}/U}\otimes\mathbb{R}$ is a constant sheaf. To be precise, this means that there exist natural isomorphisms
${\mathcal{P}}ic_{X_{U}/U}(V)\otimes\mathbb{R}\simeq{\rm{Pic}}(X_{V}/V)\simeq N^{1}(X_{V}/V)$
for any open subset $V\subset U$. Recall that by our convention, ${\rm{Pic}}(X/T)$ is the $\mathbb{R}$-vector space generated by Cartier divisors on $X$ modulo linear equivalences over $T$, which is the same as ${\rm{Pic}}(X/T)_{\mathbb{Z}}\otimes_{\mathbb{Z}}\mathbb{R}$. In fact, only certain vanishing properties are needed to ensure the statement holds. Theorem 1.3 provides a more general formulation.
We need several lemmas in the proof of Theorem 1.3.
First, recall the definition of the triviality of a map (see [Mat22, §11] for details). Consider a diagram of spaces and mappings in the analytic topology:
[…]
We say that $f$ is trivial over $Z$ if there exist spaces $X_{0}$ and $Y_{0}$, a mapping $f_{0}:X_{0}\to Y_{0}$, and homeomorphisms
$X\cong X_{0}\times Z,\quad Y\cong Y_{0}\times Z$
such that the following diagram is commutative:
[…]
We say that $f$ is locally trivial over $Z$ if, for any point $z\in Z$, the above property holds in some analytic neighborhood of $z$.
The following local triviality of the morphism is likely well-known to experts. For completeness, we explain how it can be deduced from the literature.
Lemma 4.1.
Let $h:Y\to X$ and $X\to T$ be projective fibrations of varieties. Then there exists a Zariski open subset $U\subset T$ such that $h_{U}:Y_{U}\to X_{U}$ is locally trivial over $U$ (in the analytic topology).
Proof.
After shrinking $T$, we can assume that there exists a Whitney-Thom stratification of $h$ (in the Zariski topology), which satisfies Thom’s $A_{h}$-condition (see [Hir77, Page 247, Theorem 2] or the proof of [Hir77, Page 248, Corollary 1]). Shrinking $T$ further, by Thom’s second isotopy lemma (see [Mat22, Proposition 11.1]), we can assume that the morphism $h$ is locally trivial over $T$. ∎
Lemma 4.2.
Let $g:Y\to X$ be a projective morphism between normal varieties with $R^{1}g_{*}\mathcal{O}_{Y}=0$. If $D$ is a $\mathbb{Q}$-Cartier divisor on $Y$ with $D\equiv 0/X$, then $D\sim_{\mathbb{Q}}0/X$.
In particular, suppose that $g$ is a birational morphism and that $(X,\Delta)$ is a klt pair for some effective $\mathbb{R}$-divisor $\Delta$ on $X$. If $D$ is a $\mathbb{Q}$-Cartier divisor on $Y$ with $D\equiv 0/X$, then $g_{*}D$ is a $\mathbb{Q}$-Cartier divisor on $X$ such that $g^{*}(g_{*}D)=D$.
Proof.
The argument below closely follows the proof of [KM92, Proposition (12.1.4)] and strengthens it. For the reader’s convenience, we include the full proof.
Let $x\in X$ be a point. Then there exists a triangulation of the topological space $Y$ such that the analytic subset $g^{-1}(x)$ appears as the support of a sub-complex (see [BHPV04, I, Theorem (8.8)]). Let $x\in U$ be a sufficiently small Stein neighborhood. Then $g^{-1}(U)$ deformation retracts onto the sub-complex $g^{-1}(x)$ (see [Hat00, Appendix, Proposition A.5]). In particular, we have
$H^{i}(g^{-1}(U),\mathbb{Z})=H^{i}(g^{-1}(x),\mathbb{Z}),i\in\mathbb{Z}_{\geq 0}.$
Let $(-)^{\rm an}$ denote the analytification functor. As $R^{1}g^{\rm an}_{*}\mathcal{O}^{\rm an}_{Y}=(R^{1}g_{*}\mathcal{O}_{Y})^{\rm an}=0$ (see [SGA1, Exposé XII, Théorème 4.2]) and $H^{1}(U,\mathcal{O}^{\rm an}_{X})=0$, we have
$H^{1}(g^{-1}(U),\mathcal{O}^{\rm an}_{Y})=0.$
By the exponential sequence, we have
$H^{1}(g^{-1}(U),\mathcal{O}^{\rm an}_{Y})\to H^{1}(g^{-1}(U),\mathcal{O}_{Y}^{\rm an,*})\to H^{2}(g^{-1}(U),\mathbb{Z}),$
and thus induces the natural injective group homomorphism
${{\rm{Pic}}}(g^{-1}(U))\hookrightarrow H^{2}(g^{-1}(U),\mathbb{Z})=H^{2}(g^{-1}(x),\mathbb{Z}).$
If $D\equiv 0/X$, then $D|_{g^{-1}(x)}\equiv 0$ and thus $D|_{g^{-1}(U)}$ is a torsion element of $H^{2}(g^{-1}(x),\mathbb{Z})$. If $d$ is the order of the torsion group of $H^{2}(g^{-1}(x),\mathbb{Z})$, then $dD|_{g^{-1}(U)}=0\in{\rm{Pic}}(g^{-1}(U))$. Note that there are only finitely many topological types of the fibers of $g$ (for example, using Lemma 4.1). We can take $d\in\mathbb{N}$ sufficiently divisible, such that for each $x^{\prime}\in X$, there exists an analytic neighborhood $x^{\prime}\in U^{\prime}$ such that $dD|_{g^{-1}{(U^{\prime})}}\sim 0$ on $g^{-1}{(U^{\prime})}$. As $g^{\rm an}_{*}\mathcal{O}^{\rm an}_{Y}(dD)\simeq(g_{*}\mathcal{O}_{Y}(dD))^{\rm an}$ (see [SGA1, Exposé XII, Théorème 4.2]), we have that $g_{*}\mathcal{O}_{Y}(dD)$ is a line bundle on $X$. By the natural sheaf morphism, $g^{*}g_{*}(\mathcal{O}_{Y}(dD))\to\mathcal{O}_{Y}(dD)$, we have the natural sheaf morphism (in the Zariski topology)
$\theta:\mathcal{O}_{Y}\to\mathcal{O}_{Y}(dD)\otimes(g^{*}g_{*}(\mathcal{O}_{Y}(dD)))^{\vee}.$
On $g^{-1}(U^{\prime})$, as $g^{\rm an,*}g^{\rm an}_{*}(\mathcal{O}^{\rm an}_{Y}(dD))\to\mathcal{O}^{\rm an}_{Y}(dD)$ is surjective and $\mathcal{O}^{\rm an}_{Y}$ is torsion free, we see that $\theta^{\rm an}|_{g^{-1}(U^{\prime})}$ is an isomorphism. Thus $\theta^{\rm an}$ is an isomorphism. This implies that $\theta$ is an isomorphism (for example see [SGA1, Exposé XII, Théorème et définition 1.1]). This shows $D\sim_{\mathbb{Q}}0/X$.
Next, assume that $g$ is birational and there exists a divisor $\Delta$ such that $(X,\Delta)$ has klt singularities. Replacing $Y$ by a resolution, we can assume that $Y$ is smooth. As klt singularities are rational singularities, we have $R^{1}g_{*}\mathcal{O}_{Y}=0$. Hence, we have $D\sim_{\mathbb{Q}}0/X$ by the previous discussion. This implies that $g_{*}D$ is a $\mathbb{Q}$-Cartier divisor on $X$ and $D=g^{*}(g_{*}D)$ by the negativity lemma. This completes the proof.
Finally, we include an alternative proof of this special case below, which is more algebraic in nature.
Let $h:\tilde{X}\to X$ be a small $\mathbb{Q}$-factorial modification. Then we have
$K_{\tilde{X}}+\tilde{\Delta}=h^{*}(K_{X}+\Delta),$
where $\tilde{\Delta}$ is the strict transform of $\Delta$. Replacing $Y$ by a higher model, we can assume that $g$ factors through $h$, and let $u:Y\to\tilde{X}$ be the corresponding morphism. We have $D\equiv 0/\tilde{X}$. Thus $\tilde{B}\coloneqq u_{*}D$ is a $\mathbb{Q}$-Cartier divisor as $\tilde{X}$ is $\mathbb{Q}$-factorial and $D=u^{*}(u_{*}D)$ by the negativity lemma. In particular, we have $\tilde{B}\equiv 0/X$. By the base-point free theorem, $\tilde{B}$ is semi-ample over $X$. Hence $g_{*}D=h_{*}\tilde{B}$ is a $\mathbb{Q}$-Cartier divisor on $X$ and $h^{*}(h_{*}\tilde{B})=\tilde{B}$. Thus, we have $g^{*}(g_{*}D)=u^{*}(h^{*}(h_{*}\tilde{B}))=D$. ∎
Remark 4.3.
Lemma 4.2 does not hold for an arbitrary birational contraction $g$. Assume that $X$ is a normal projective surface which is not $\mathbb{Q}$-factorial. Let $H$ be a Weil divisor on $X$ which is not $\mathbb{Q}$-Cartier. For a resolution $g:Y\to X$, one can define the Mumford pull-back of $H$, denoted by $g^{*}H$. Then we have $g^{*}H\equiv 0/X$ by construction. However, we do not have $g^{*}H\sim_{\mathbb{Q}}0/X$, since this would imply that $H$ is $\mathbb{Q}$-Cartier. Lemma 4.2 is consistent with the fact that for projective surfaces, rational singularities are $\mathbb{Q}$-factorial (see [Bad01, Theorem 4.6]).
Proof of Theorem 1.3.
We first prove Theorem 1.3 (1), and then show that Theorem 1.3 (2) can be deduced from Theorem 1.3 (1). From now on, we assume that the natural restriction map
$N^{1}(X/T)\to N^{1}(X_{t}),\quad[D]\mapsto[D|_{X_{t}}]$ (4.2)
is surjective for very general $t\in T$.
Shrinking $T$, by [Har77, Chapter III, Theorem 12.8, Corollary 12.9], we can assume that
$R^{1}f_{*}\mathcal{O}_{X}=R^{2}f_{*}\mathcal{O}_{X}=0\text{~and~}H^{1}(X_{t},\mathcal{O}_{X_{t}})=H^{2}(X_{t},\mathcal{O}_{X_{t}})=0~\forall~t\in T.$ (4.3)
Shrinking $T$ further, we can assume that $T$ is smooth. Then, we have
${\rm{Pic}}(X_{U}/U)=N^{1}(X_{U}/U)\text{~for any open~}U\subset T\text{~and~}{\rm{Pic}}(X_{t})=N^{1}(X_{t})~\forall~t\in T.$ (4.4)
Indeed, ${\rm{Pic}}(X_{t})=N^{1}(X_{t})$ follows from $H^{1}(X_{t},\mathcal{O}_{X_{t}})=0$. We explain that ${\rm{Pic}}(X_{U}/U)=N^{1}(X_{U}/U)$. Let $\eta\in U$ be the generic point. Then, by the flat base change, we have $H^{1}(X_{\eta},\mathcal{O}_{X_{\eta}})=0$. In particular, we have ${\rm{Pic}}(X_{\eta})=N^{1}(X_{\eta})$. Thus, if $D$ is a Cartier divisor such that $D\equiv 0/U$, then we have $D_{\eta}\sim_{\mathbb{Q}}0$. Replacing $D$ by a multiple, we can assume that $D_{\eta}\sim 0$, and thus there exists $\alpha\in K(X_{\eta})=K(X_{U})$ such that $D_{\eta}={\rm div}(\alpha)$ on $X_{\eta}$. This implies that $D-{\rm div}(\alpha)$ is a vertical divisor on $X_{U}/U$. As $D-{\rm div}(\alpha)\equiv 0/U$ and $U$ is smooth, there exists a divisor $L$ on $U$ such that $D-{\rm div}(\alpha)=f^{*}L$ by the negativity lemma. That is, $D\sim_{\mathbb{Q}}0/U$. This shows ${\rm{Pic}}(X_{U}/U)=N^{1}(X_{U}/U)$.
Let $\mathcal{B}_{i},1\leq i\leq l$ be Cartier divisors on $X$ such that $\{[\mathcal{B}_{i}]\mid 1\leq i\leq l\}$ spans the vector space $N^{1}(X/T)$. Shrinking $T$, we can assume that each irreducible component of $\mathcal{B}_{i}$ is flat over $T$. By (4.2), we can further assume that for very general $t\in T$, any divisor $B_{t}$ of $N^{1}(X_{t})$ is a restriction of a divisor $\mathcal{B}$ of $N^{1}(X/T)$, and each irreducible component of $\mathcal{B}$ is flat over $T$.
By the discussion preceding this proposition, we have
${\rm{Pic}}(X_{U}/U)_{\mathbb{Z}}\subset{\mathcal{P}}ic_{X_{T}/T}(U)$
for any open subset $U\subset T$. Taking direct image sheaves of the short exact sequence,
$0\to\mathbb{Z}\to\mathcal{O}_{X_{T}}\to\mathcal{O}^{*}_{X_{T}}\to 0,$
we have the long exact sequence
$\cdots\to R^{1}f_{*}\mathcal{O}_{X_{T}}\to R^{1}f_{*}\mathcal{O}^{*}_{X_{T}}\to R^{2}f_{*}\mathbb{Z}\to R^{2}f_{*}\mathcal{O}_{X_{T}}\to\cdots.$
By (4.3), we have $R^{1}f_{*}\mathcal{O}_{X_{T}}=R^{2}f_{*}\mathcal{O}_{X_{T}}=0$. Therefore, we have $R^{1}f_{*}\mathcal{O}^{*}_{X_{T}}\simeq R^{2}f_{*}\mathbb{Z}$, and thus
${\mathcal{P}}ic_{X_{T}/T}(U)=H^{0}(U,R^{1}f_{*}\mathcal{O}^{*}_{X_{T}})=H^{0}(U,R^{2}f_{*}\mathbb{Z})$
for any open subset $U\subset T$ by equation (4.1).
By the Thom-Whitney stratification and Thom’s first isotopy lemma, there exists a constructible stratification in the Zariski topology such that each stratum is locally trivial in the analytic topology (see [ZS10, Theorem 3.2]). In particular, $R^{2}f_{*}\mathbb{Z}$ restricting to each stratum is a local system in the analytic topology (see [ZS10, Proposition 3.5]). That is, $R^{2}f_{*}\mathbb{Z}$ restricting to each stratum is a locally constant sheaf in the analytic topology. Shrinking $T$, we can assume that $R^{2}f_{*}\mathbb{Z}$ is a local system over $T$.
The key of the argument below is to show that, in the above setting, $R^{2}f_{*}\mathbb{R}=(R^{2}f_{*}\mathbb{Z})\otimes\mathbb{R}$ is indeed a constant sheaf on $T$ in the Zariski topology. The argument in this part is similar to the argument of [dFH11, Lemma 6.6].
Note that the above discussion gives the relations
$N^{1}(X_{U}/U)={\rm{Pic}}(X_{U}/U)\subset{\mathcal{P}}ic_{X_{T}/T}(U)\otimes\mathbb{R}=H^{0}(U,R^{2}f_{*}\mathbb{R})\subset H^{0}(\mathbb{D},R^{2}f_{*}\mathbb{R})$ (4.5)
for any open subset $U\subset T$ and any analytic open subset $\mathbb{D}\subset U$. We will show that these inclusions are indeed equalities. By assumption (see (4.2)), the restriction map
${\rm{Pic}}(X_{U}/U)\to{\rm{Pic}}(X_{t})$
is surjective for very general $t\in U$. Let $t_{0}\in T$ be an arbitrary point. As $R^{2}f_{*}\mathbb{Z}$ is a local system, there exists a contractible analytic open subset $t_{0}\in\mathbb{D}\subset T$ such that
$H^{0}(\mathbb{D},R^{2}f_{*}\mathbb{Z})\simeq H^{2}(X_{t_{0}},\mathbb{Z}).$ (4.6)
By (4.3) and (4.4), we have ${\rm{Pic}}(X_{t})_{\mathbb{Z}}=H^{1}(X_{t},\mathcal{O}_{X_{t}}^{*})=H^{2}(X_{t},\mathbb{Z})$. Therefore, for any Cartier divisor $B_{t_{0}}\in{\rm{Pic}}(X_{t_{0}})$, there exists a Cartier divisor $B_{t}\in{\rm{Pic}}(X_{t})$ for any $t\in\mathbb{D}$ through the natural identification (4.6). By the density of very general points in analytic topology, we can take $t$ to be a very general point such that ${\rm{Pic}}(X_{U}/U)|_{X_{t}}\simeq{\rm{Pic}}(X_{t})$. Therefore, there exists a $\mathbb{Q}$-Cartier divisor $\mathcal{B}$ on $X_{U}$, which is a flat cycle over $U$, such that $\mathcal{B}_{t}\coloneqq\mathcal{B}|_{X_{t}}=B_{t}$. Moreover, as $\mathcal{B}$ is a flat cycle over $U$, $\mathcal{B}_{t_{0}}$ can be identified with $\mathcal{B}_{t}=B_{t}$ as cycles through the isomorphism
$H^{2}(X_{t_{0}},\mathbb{Z})\simeq H^{0}(\mathbb{D},R^{2}\pi_{*}\mathbb{Z})\simeq H^{2}(X_{t},\mathbb{Z}).$ (4.7)
Hence, we have $\mathcal{B}_{t_{0}}=B_{t_{0}}$ as cycles by the choice of $B_{t}$. We conclude that $\mathcal{B}_{t_{0}}=B_{t_{0}}$ as divisors by ${\rm{Pic}}(X_{t})_{\mathbb{Z}}=H^{2}(X_{t},\mathbb{Z})$. From this, we see that (4.5) are indeed equalities. Moreover, as $t_{0}$ is an arbitrary point, the above discussion also shows that the restriction map
$N^{1}(X_{U}/U)\to N^{1}(X_{t})$
is surjective for any $t\in U$.
Next, we deduce Theorem 1.3 (2) from Theorem 1.3 (1). We proceed with the argument in two steps.
Step 1. We prove the claim under the additional assumption that $X$ is smooth. We claim that there exists a generically finite base change $T^{\prime}\to T$ and an open subset $T_{0}\subset T^{\prime}$ such that
$N^{1}(X_{T_{0}}/T_{0})\to N^{1}(X_{T_{0},t})$
for very general $t\in T_{0}$. This can be proved by the same argument as [KM92, Proposition 12.2.5]. Indeed, let $\mathcal{H}$ be the Hilbert scheme that parametrizes (n1)(n-1)-cycles of fibers of $f$, where $\dim(X/T)=n$. let $\mathcal{H}=\cup_{i\in J}\mathcal{H}_{i}$ be the decomposition into irreducible components. Then $J$ is at most a countable set and the natural map $p_{i}:\mathcal{H}_{i}\to T$ is projective. Let $I=\{i\in J\mid p_{i}(\mathcal{H}_{i})\neq T\}$. Then
$W=T-\cup_{i\in I}p_{i}(\mathcal{H}_{i})$
consists of very general points.
Shrinking $T$, we can assume that $X\to T$ is a locally trivial family in the analytic topology. As $X$ is smooth, shrinking $T$ further, we can assume that $f$ is a smooth morphism. Hence, $h^{1,1}(X_{t},\mathbb{Q})$ is invariant for $t\in T$. Fix some $t_{0}\in W$. By the choice of $W$, for each prime divisor $D$ on $X_{t_{0}}$, there exists some $i\in I$ and a Weil divisor $\mathcal{D}$ on $X_{\mathcal{H}_{i}}\coloneqq\mathcal{H}_{i}\times_{T}X$ such that $\mathcal{D}|_{X_{\mathcal{H}_{i},s_{0}}}=D$, where $s_{0}\in\mathcal{H}_{i}$ is a closed point such that $p_{i}(s_{0})=t_{0}$. Note that we have $X_{\mathcal{H}_{i},s_{0}}\simeq X_{t_{0}}$. Let $\mathcal{G}_{i}\subset\mathcal{H}_{i}$ be a closed subvariety such that $\mathcal{G}_{i}\to T$ is a generically finite morphism. Then we still have $\mathcal{D}|_{X_{\mathcal{G}_{i},s_{0}}}=D$, where $s_{0}\in\mathcal{G}_{i}$ is a closed point such that $p_{i}(s_{0})=t_{0}$. As $f$ is smooth, $f_{\mathcal{G}_{i}}$ is still a smooth morphism. Shrinking $\mathcal{G}_{i}$, we can assume that $\mathcal{G}_{i}$ is a smooth variety, and thus $X_{\mathcal{G}_{i}}$ is a smooth variety. Hence $\mathcal{D}|_{\mathcal{G}_{i}}$ is Cartier divisor on $X_{\mathcal{G}_{i}}$.
As $h^{1,1}(X_{t_{0}},\mathbb{Q})$ is of finite dimension, we can repeat the above process finitely many times, and obtain a generically finite morphism $q:T^{\prime}\to T$ and an open subset $T_{0}\subset T^{\prime}$ such that
$N^{1}(X_{T_{0}}/T_{0})\to N^{1}(X_{T_{0},t_{0}})$
is surjective. Moreover, each Cartier divisor $B$ on $X_{T_{0},t_{0}}$ is a restriction of a cycle $\mathcal{B}$ on $X_{T_{0}}$ which is flat over $T_{0}$. Note that $X\to T$ is a locally trivial family in the analytic topology. Suppose that $t_{0}\in\mathbb{D}\subset T_{0}$ is a contractible analytic open set such that $X_{\mathbb{D}}$ is trivial. By the flatness of $\mathcal{B}$ over $T_{0}$, for each $\in\mathbb{D}, [t][\mathcal{B}_{t}$] corresponds to $\mathcal{B}_{t_{0}}]$ under the identification $H^{2}(X_{t_{0}},\mathbb{Q})\simeq H^{2}(X_{t},\mathbb{Q})$. By the connectedness of $T_{0}$, we see that
$N^{1}(X_{T_{0}}/T_{0})\to N^{1}(X_{T_{0},t})$
is surjective for each $t\in q^{-1}(W)\cap T_{0}$. This verifies the conditions in Theorem 1.3 (1). Note that for the dense subset $S^{\prime}=q^{-1}(S)$, each point $s\in S^{\prime}$ satisfies that the fiber $X_{T_{0},s}$ has rational singularities and $H^{i}(X_{T_{0},s},\mathcal{O}_{X_{T_{0},s}})=0,i=1,2$. The above shows that Theorem 1.3 (2) holds when $X$ is smooth.
Step 2. We show the claim in the general case.
Let $h:Y\to X$ be a resolution which can be assumed to be a fiber-wise resolution after shrinking $T$. Note that for the dense subset $S$, each point $s\in S$ satisfies that the fiber $Y_{s}$ has rational singularities and $H^{i}(Y_{s},\mathcal{O}_{Y_{s}})=0,i=1,2$. By Step 1, there exist a generically finite base change $q:T^{\prime}\to T$ and an open subset $T_{0}\subset T^{\prime}$, such that $N^{1}(Y_{T_{0}}/T_{0})\to N^{1}(Y_{T_{0},t})$ is surjective for $t\in q^{-1}(W)\cap T_{0}$, where $W\subset T$ consists of very general points. Moreover, each Cartier divisor $B$ on $Y_{T_{0},t_{0}}$ is a restriction of a cycle $\mathcal{B}$ on $Y_{T_{0}}$ which is flat over $T_{0}$.
The above morphisms are summarized in the following diagram:
[…]
As before, after shrinking $T$, we can assume that $f$ is locally trivial in the analytic topology. Moreover, by Lemma 4.1, the morphism $h$ can also be locally trivialized over $T$ in the analytic topology. In particular, this implies that for a fixed $t_{0}\in T$, there exists an analytic open subset $t_{0}\in\mathbb{D}\subset T$, such that for any $t\in\mathbb{D}$, the following diagram commutes
[…]
where $H^{2}(Y_{t_{0}},\mathbb{Z})\simeq H^{2}(Y_{t},\mathbb{Z})$ and $H^{2}(X_{t_{0}},\mathbb{Z})\simeq H^{2}(X_{t},\mathbb{Z})$ are obtained through natural identifications as (4.7).
Take any $t_{0}\in W\cap q(T_{0})$. Note that by (4.3) and (4.4), $\dim N^{1}(X_{t})$ is invariant for each $t\in T$. Let
$h^{*}(N^{1}(X_{t_{0}}))\coloneqq\{[h^{*}B]\mid[B]\in N^{1}(X_{t_{0}})\}$
be the subspace of $N^{1}(Y_{t_{0}})$. Let $s_{0}\in q^{-1}(W)\cap T_{0}$ be a closed point such that $q(s_{0})=t_{0}$. By the previous discussion, there are finitely many cycles $\mathcal{B}_{i},1\leq i\leq l$ on $Y_{T_{0}}$ which are flat over $T_{0}$ such that $\{[\mathcal{B}_{i,s_{0}}]\in N^{1}(Y_{T_{0},s_{0}})\}$ generates the subspace $h^{*}(N^{1}(X_{t_{0}}))$.
In the analytic topology, as $f$ is locally trivial and $h$ is locally trivial over $T$, we have that $f\circ h:Y\to T$ is also locally trivial. The corresponding property is preserved after the base change $T^{\prime}\to T$. In particular, there exists an analytic open subset $s_{0}\in\mathbb{D}\subset T_{0}$ such that for any $s\in\mathbb{D}$, the following diagram commutes
[…]
where $H^{2}(Y_{T_{0},s_{0}},\mathbb{Z})\simeq H^{2}(Y_{T_{0},s},\mathbb{Z})$ and $H^{2}(X_{T_{0},s_{0}},\mathbb{Z})\simeq H^{2}(X_{T_{0},s},\mathbb{Z})$ are obtained through natural identifications as (4.7). As $\mathcal{B}_{i,s_{0}}=h_{T_{0},s_{0}}^{*}B$ for some $\mathbb{Q}$-Cartier divisor $B$ on $X_{T_{0},s_{0}}$ by construction, we have $\mathcal{B}_{i,s_{0}}\equiv 0/X_{T_{0},s_{0}}$. Thus, we have $\mathcal{B}_{i,s}\equiv 0/X_{T_{0},s}$ for $s\in\mathbb{D}$ by (4.8). By the connectedness of $T_{0}$, we have
$\mathcal{B}_{i}\equiv 0/X_{T_{0}}.$
As general fibers of $f_{T_{0}}$ have rational singularities, after shrinking $T_{0}$, we can assume that $X_{T_{0}}$ still have rational singularities (see [Elk78]). By Lemma 4.2, we have $\mathcal{B}_{i}\sim_{\mathbb{Q}}0/X_{T_{0}}$. As $h_{T_{0}}$ is a birational morphism, $\mathcal{D}_{i}\coloneqq h_{T_{0},*}\mathcal{B}_{i}$ is a $\mathbb{Q}$-Cartier divisor and $\mathcal{B}_{i}=h_{T_{0}}^{*}\mathcal{D}_{i}$ by the negativity lemma. Therefore, we have
$h^{*}_{T_{0},s_{0}}(\mathcal{D}_{i,s_{0}})\sim_{\mathbb{Q}}h^{*}_{T_{0},s_{0}}(B),$
and hence $\mathcal{D}_{i,s_{0}}\sim_{\mathbb{Q}}B$. This shows that
$N^{1}(X_{T_{0}}/T_{0})\to N^{1}(X_{T_{0},s_{0}})$
is surjective.
Finally, because $\mathcal{B}_{i},1\leq i\leq l$ are cycles on $Y_{T_{0}}$ which are flat over $T_{0}$, and $Y_{T_{0}}$ is locally trivial over $T_{0}$ in the analytic topology, we see that $\{[\mathcal{B}_{i,s}]\mid 1\leq i\leq l\}$ spans a subspace of the same dimension as $\dim N^{1}(X_{t_{0}})$. This implies that $\{[\mathcal{D}_{i,s}]\mid 1\leq i\leq l\}$ spans a subspace of the same dimension as $\dim N^{1}(X_{t_{0}})$. Therefore, the natural map
$N^{1}(X_{T_{0}}/T_{0})\to N^{1}(X_{T_{0},s})$
is surjective for each $s\in q^{-1}(W)\cap T_{0}$. This verifies that the conditions of Theorem 1.3 (1) are satisfied, and thus the desired claim follows. ∎
Corollary 4.4.
Let $f:X\to T$ be a projective fibration. Suppose that $S\subset T$ is a Zariski dense subset such that for any $s\in S$, the fiber $X_{s}$ is a rationally chain-connected variety with rational singularities. Then, up to a generically finite base change of $T$, there exists a non-empty open subset $T_{0}\subset T$, such that ${\mathcal{P}}ic_{X_{T_{0}}/T_{0}}\otimes\mathbb{R}$ is a constant sheaf. Moreover, for any open subset $U\subset T_{0}$, the natural restriction maps
$N^{1}(X_{T_{0}}/T_{0})\to N^{1}(X_{U}/U)\to N^{1}(X_{t})$
are isomorphisms for any $t\in U$. In particular, if $X_{s},s\in S$ are varieties of Fano type, then the above claim holds.
Proof.
Shrinking $T$, we have $H^{1}(X_{s},\mathcal{O}_{X_{s}})=H^{2}(X_{s},\mathcal{O}_{X_{s}})=0$ for each $s\in S$ as $X_{s}$ is a rationally chain-connected variety with rational singularities. Hence, the claim follows from Theorem 1.3. Besides, a Fano type variety is naturally a rationally chain-connected variety with rational singularities. ∎
4.2. Deformation of nef cones
Theorem 4.5.
Let $f:X\to T$ be a projective fibration. Assume that ${\rm{Nef}}(X_{t})$ is a rational polyhedral cone for each $t\in T$. Suppose that for any open subset $V\subset T$,
  1. (1)
    the natural restriction map $N^{1}(X_{V}/V)\to N^{1}(X_{t}),t\in V$ is an isomorphism, and
  2. (2)
    for any $\mathcal{D}\in\overline{\rm{Eff}}(X_{V}/V)$, there exsits a $\mathcal{D}$-MMP/V/V which terminates to a model such that the push-foreard of $\mathcal{D}$ is nef.
Then there exists a non-empty Zariski open subset $T_{0}\subset T$, such that for any open subset $U\subset T_{0}$, we have
${\rm{Nef}}(X_{T_{0}}/T_{0})\simeq{\rm{Nef}}(X_{U}/U)\simeq{\rm{Nef}}(X_{t}),t\in U$
under the natural restriction maps. Moreover, there are isomorphisms of Mori cones
${\rm{NE}}(X_{t})\simeq{\rm{NE}}(X_{U}/U)\simeq{\rm{NE}}(X_{T_{0}}/T_{0}),t\in U$
under the natural inclusion maps.
Proof.
If we have $\mathcal{D}\in N^{1}(X/T)$ such that $\mathcal{D}_{t}\coloneqq\mathcal{D}|_{X_{t}}$ is nef, then $\mathcal{D}_{t}$ is nef for very general $t\in T$ by the openness of ampleness. Set
[…]
We claim that $\mathcal{N}$ is an open subset. By assumption, we can run some $\mathcal{D}$-MMP/$T$, $X\dasharrow X^{\prime}$, such that $\mathcal{D}^{\prime}$ is nef/T/T, where $\mathcal{D}^{\prime}$ is the strict transform of $\mathcal{D}$ on $X^{\prime}$. This MMP is invariant when restricting to each fiber $X_{t},t\in\mathcal{N}_{\mathcal{D}}$. As the image $Z\subset T$ of the contracting/flipping loci is a closed subset of $T$, and since $Z=T\setminus\mathcal{N}_{\mathcal{D}}$, it follows that $\mathcal{N}_{\mathcal{D}}$ is an open subset of $T$.
Note that ${\rm{Nef}}(X_{t})$ is a rational polyhedral cone by assumption. Suppose that ${\rm{Nef}}(X_{t})$ is generated by finite rays $\mathcal{D}_{i},i=1,\ldots,l$. Then, the fact that $\mathcal{N}_{\mathcal{D}_{i}}$ is open for each $i$ implies that the map $t\mapsto{\rm{Nef}}(X_{t})$ is lower semi-continuous in the following sense: for any $t_{0}\in T$, there exists a Zariski open set $U_{t_{0}}$ containing $t_{0}$ such that ${\rm{Nef}}(X_{t})$ contains ${\rm{Nef}}(X_{t_{0}})$ for any $t\in U_{t_{0}}$. Here, we view ${\rm{Nef}}(X_{t})$ as a subset of $N^{1}(X/T)\cong N^{1}(X_{t})$ by assumption, and the cones are compared under this natural identification. By abuse of notation, the inclusion $\subset$” is understood under this identification.
Next, we show that there exists an open subset $T_{0}\subset T$ such that ${\rm{Nef}}(X_{t})$ is invariant for all $t\in T_{0}$. We argue by induction on the dimension of the base $T$. The case when $\dim T=0$ is obvious. Assume that the claim holds when $\dim T=n-1$. Then we show that it also holds when $\dim T=n$.
First, we choose countably many hypersurfaces $\{Z_{i}\}$ in $T$ such that $\cup Z_{i}$ is Zariski dense in $T$. For each $Z_{i}$, by induction, there exists an open subset $U_{i}\subset T$ such that $U_{i}\cap Z_{i}$ is non-empty and ${\rm{Nef}}(X_{t})$ is invariant on $U_{i}\cap Z_{i}$. We set $P_{i}\coloneqq{\rm{Nef}}(X_{t_{i}})$ for some $t_{i}\in U_{i}\cap Z_{i}$. By the lower semi-continuity, after shrinking $U_{i}$ around $t_{i}$, we can assume that $P_{i}\subset{\rm{Nef}}(X_{t})$ for any $t\in U_{i}$. By abuse of notation, we also write $P_{i}$ for the cone inside $N^{1}(X/T)$ under the natural identification $N^{1}(X/T)\simeq N^{1}(X_{t_{i}})$. Let
$P_{\infty}\coloneqq\overline{\cup_{i}P_{i}}$
be the closure of $\cup_{i}P_{i}$ inside $N^{1}(X/T)$. Set $U_{\infty}\coloneqq\cap_{i}U_{i}$. We claim that $P_{\infty}={\rm{Nef}}(X_{t})$ for any $t\in U_{\infty}$. First, we have $P_{\infty}\subset{\rm{Nef}}(X_{t})$ for any $t\in U_{\infty}$ as $P_{i}\subset{\rm{Nef}}(X_{t})$ for $t\in U_{i}$. Suppose that $P_{\infty}\subsetneqq{\rm{Nef}}(X_{t_{0}})$ for some $t_{0}\in U_{\infty}$, then, by the lower semi-continuity, there exists an open subset $t_{0}\in V$ such that $P_{\infty}\subsetneqq{\rm{Nef}}(X_{t})$ for any $t\in V$. Since $\cup_{i}Z_{i}$ is dense in $T$, $V$ must intersect some $Z_{i}\cap U_{i}$. Thus, we have
$P_{\infty}\subsetneqq{\rm{Nef}}(X_{t})$
for some $t\in U_{i}\cap V$. However, we have ${\rm{Nef}}(X_{t})=P_{i}$ for any $t\in U_{i}\cap Z_{i}$ by construction. This is a contradiction. Therefore, we have $P_{\infty}={\rm{Nef}}(X_{t})$ for any $t\in U_{\infty}$.
For any $t^{\prime}\in U_{\infty}$, there exists an open subset $t^{\prime}\in V_{0}$ such that $P_{\infty}={\rm{Nef}}(X_{t^{\prime}})\subset{\rm{Nef}}(X_{t})$ for each $t\in V_{0}$. If there exists some $\tilde{t}$ such that the inclusion is strict, then, by the lower semi-continuity, there exists an open subset $\tilde{t}\in\tilde{V}_{0}$ such that $P_{\infty}\subsetneqq{\rm{Nef}}(X_{t})$ for any $t\in\tilde{V}_{0}$. However, as $U_{\infty}$ consists of very general points, we have $U_{\infty}\cap\tilde{V}_{0}\neq\emptyset$. This is a contradiction as ${\rm{Nef}}(X_{t^{\prime\prime}})=P_{\infty}$ for any $t^{\prime\prime}\in U_{\infty}$ as shown above. This shows that $P_{\infty}={\rm{Nef}}(X_{t})$ for each $t\in V_{0}$. Therefore, the natural restriction map
${\rm{Nef}}(X_{V_{0}}/V_{0})\to{\rm{Nef}}(X_{t}),t\in V_{0}$
is an isomorphism. In fact, if $\mathcal{D}\in N^{1}(X_{V_{0}}/V_{0})$ is a divisor such that $\mathcal{D}_{t}\in{\rm{Nef}}(X_{t})$. Then $\mathcal{D}_{t^{\prime}}\in N^{1}(X_{t^{\prime}})$ is the divisor identified with $\mathcal{D}_{t}$ for each $t^{\prime}\in V_{0}$. As ${\rm{Nef}}(X_{t})=P_{\infty}={\rm{Nef}}(X_{t^{\prime}})$ for $t,t^{\prime}\in V_{0}$, we see that $\mathcal{D}_{t^{\prime}}\in{\rm{Nef}}(X_{t^{\prime}})$. In particular, we have $\mathcal{D}\in{\rm{Nef}}(X_{V_{0}}/V_{0})$.
Denote $T_{0}:=V_{0}$. For any open subset $U\subset T_{0}$, we have the natural inclusions
${\rm{Nef}}(X_{T_{0}}/T_{0})|_{X_{t}}\hookrightarrow{\rm{Nef}}(X_{U}/U)|_{X_{t}}\hookrightarrow{\rm{Nef}}(X_{t})$
for any $t\in U$. As ${\rm{Nef}}(X_{T_{0}}/T_{0})|_{X_{t}}={\rm{Nef}}(X_{t})$, we have ${\rm{Nef}}(X_{U}/U)|_{X_{t}}={\rm{Nef}}(X_{t})$. As the natural restriction map $N^{1}(X_{U}/U)\to N^{1}(X_{t})$ is an isomorphism, we have ${\rm{Nef}}(X_{U}/U)\simeq{\rm{Nef}}(X_{t})$ for any $t\in U$.
For the last statement on the isomorphism of Mori cones, the natural maps
${\rm{NE}}(X_{t})\to{\rm{NE}}(X_{U}/U)\to{\rm{NE}}(X_{T_{0}}/T_{0}),t\in U$
are injective as $N^{1}(X_{T_{0}}/T_{0})\simeq N^{1}(X_{U}/U)\simeq N^{1}(X_{t})$. Thus, we have ${\rm{NE}}(X_{t})\subset{\rm{NE}}(X_{U}/U)$ under this identification. If the inclusion is strict, then there exists a divisor $\mathcal{D}$ on $X_{U}/U$ such that $\mathcal{D}$ is not nef/U/U while $\mathcal{D}_{t}\in{\rm{Nef}}(X_{t})$. By ${\rm{Nef}}(X_{U}/U)\simeq{\rm{Nef}}(X_{t})$, we see that $\mathcal{D}$ is nef/U/U. This is a contradiction. The remaining isomorphisms can be shown similarly. ∎
As a corollary of Theorem 4.5, we have the following two corollaries.
Corollary 4.6.
Let $f:X\to T$ be a projective fibration. Assume that $X_{t}$ is a Mori dream space for each $t\in T$. Suppose that for any open subset $V\subset T$,
  1. (1)
    the natural restriction map $N^{1}(X_{V}/V)\to N^{1}(X_{t}),t\in V$ is an isomorphism, and
  2. (2)
    $X_{V}$ is a Mori dream space over $V$.
Then there exists a non-empty Zariski open subset $T_{0}\subset T$, such that for any open subset $U\subset T_{0}$, we have
${\rm{Nef}}(X_{T_{0}}/T_{0})\simeq{\rm{Nef}}(X_{U}/U)\simeq{\rm{Nef}}(X_{t}),t\in U$
under the natural restriction maps. Moreover, there are isomorphisms of Mori cones
${\rm{NE}}(X_{t})\simeq{\rm{NE}}(X_{U}/U)\simeq{\rm{NE}}(X_{T_{0}}/T_{0}),t\in U$
under the natural inclusion maps.
Proof.
As $X_{t}$ is a Mori dream space, ${\rm{Nef}}(X_{t})$ is a rational polyhedral cone. As $X_{V}$ is a Mori dream space over $V$, for any pseudo-effective divisor $\mathcal{D}$ on $X_{V}$ over $V$, there exists a $\mathcal{D}$-MMP/V/V that terminates to a model such that the push-forward of $\mathcal{D}$ is semi-ample. This verifies the conditions in Theorem 4.5, and thus the claim follows. ∎
Corollary 4.7.
Let $f:X\to T$ be a projective fibration. Suppose that there exists a Zariski dense subset $S\subset T$ such that the fibers $X_{s}$ for $s\in S$ satisfy one of the cases in Theorem 1.2. Then, up to a generically finite base change of $T$, there exists a non-empty Zariski open subset $T_{0}\subset T$, such that for any open subset $U\subset T_{0}$, we have
${\rm{Nef}}(X_{T_{0}}/T_{0})\simeq{\rm{Nef}}(X_{U}/U)\simeq{\rm{Nef}}(X_{t}),t\in U$
under the natural restriction maps. Moreover, there are isomorphisms of Mori cones
${\rm{NE}}(X_{t})\simeq{\rm{NE}}(X_{U}/U)\simeq{\rm{NE}}(X_{T_{0}}/T_{0}),t\in U$
under the natural inclusion maps.
Proof.
By Theorem 1.2, we can assume that $X$ is of Fano type over $T$ after shrinking $T$. Shrinking $T$ further, we can assume that $X_{t}$ is of Fano type for each $t\in T$. Note that a variety of Fano type (resp. over $T$) is a Mori dream space (resp. over $T$). Taking a generically finite base change of $T$ and shrinking $T$ again, we may assume that Corollary 4.6 (1) is satisfied by Theorem 1.3 (2). Therefore, the claim follows from Corollary 4.6. ∎
5. MMP in a family and deformations of effective cones and Mori chamber decompositions
This section is devoted to the proof of Theorem 1.4, Theorem 1.6, as well as the invariance of effective cones and Mori chamber decompositions stated in Theorem 1.5.
5.1. Stable boundedness
Let $X$ be of Fano type over $T$. Set
[…]
Note that the above set includes not only varieties birationally contracted by $X$ but also the birational contraction maps from $X$. However, for the sake of simplicity, we also write $Y/T\in{\rm{bc}}(X/T)$ instead of $h:X\dasharrow Y/T]\in{\rm{bc}}(X/T)$. For $Y_{1}/T,Y_{2}/T\in{\rm{bc}}(X/T)$, $Y_{1}\simeq Y_{2}/T$ should be understood as that there exists an isomorphism $\theta:Y_{1}\to Y_{2}$, such that $h_{2}=\theta\circ h_{1}:X\dasharrow Y_{2}$, where $h_{i}:X\dasharrow Y_{i},i=1,2$ are given birational contractions. Moreover, for a subvariety $V\subset T$, we write $Y_{V}/V$ instead of the base change of $h$ to $V$.
To study an MMP of a Fano type variety $X/T$, we also consider a subset ${\rm{bc}}(X/T)_{\rm MMP}$ of ${\rm{bc}}(X/T)$ which we define below. First, we call a sequence of birational contractions
$X=X_{0}\dashrightarrow X_{1}\dashrightarrow\cdots\dashrightarrow X_{n}/T$
a partial MMP/T/T (with respect to a divisor $D$) if each step consists of $D_{i}$-extremal divisorial/flipping contractions and $D_{i}$-flips, where $D_{i}$ is the strict transform of $D$ on $X_{i}$. Note that $D_{n}$ is not required to be nef over $T$. On the other hand, if $D_{n}$ becomes nef and big over $T$, and $X_{n}\to X_{n+1}$ is the birational contraction induced by $D_{n}$, then
$X=X_{0}\dashrightarrow X_{1}\dashrightarrow\cdots\dashrightarrow X_{n}\to X_{n+1}/T$
is also called a partial MMP/T/T. We set
[…]
Later, we will show that ${\rm{bc}}(X/T)_{\rm MMP}={\rm{bc}}(X/T)$ when $X$ is $\mathbb{Q}$-factorial in Proposition 5.4.
Let $\tilde{X}\to X$ be a small $\mathbb{Q}$-factorial modification. Then ${\rm{bc}}(\tilde{X}/T)$ can be naturally identified with ${\rm{bc}}(X/T)$. By abuse of notation, we will write ${\rm{bc}}(\tilde{X}/T)={\rm{bc}}(X/T)$ for this identification. The following is a special case of [BCHM10, Corollary 1.1.5].
Proposition 5.1.
Let $X$ be of Fano type over $T$. Then ${\rm{bc}}(X/T)$ is a finite set.
In fact, we have a stronger statement that shows the birational contractions eventually stabilize in the process of shrinking $T$.
Proposition 5.2.
Let $X$ be of Fano type over $T$. Then there exists an open subset $T_{0}\subset T$ such that for any open subset $V\subset T_{0}$ and $Y/V\in{\rm{bc}}(X_{V}/V)$, there exists an element $W/T_{0}\in{\rm{bc}}(X_{T_{0}}/T_{0})$ such that $W_{V}/V$ is isomorphic to $Y/V$.
Proof.
We proceed with the argument in several steps.
Step 1. Let $K\coloneqq K(T)$ be the function field of $T$ and $\bar{K}$ the algebraic closure. By Proposition 5.1, the set ${\rm{bc}}(X_{\bar{K}}/\bar{K})$ is finite. Suppose that we have ${\rm{bc}}(X_{\bar{K}}/\bar{K})=\{W_{1},\cdots,W_{l}\}$. Then for any open subset $V\subset T$, and $Y/V\in{\rm{bc}}(X_{V}/V)$, we have $Y_{\bar{K}}\simeq W_{i}\in{\rm{bc}}(X_{\bar{K}}/\bar{K})$ for some $i$. For each $i$, if there exist an open subset $V\subset T$ and $Y/V\in{\rm{bc}}(X_{V}/V)$ such that $Y_{\bar{K}}\simeq W_{i}\in{\rm{bc}}(X_{\bar{K}}/\bar{K})$, then we fix one such $V$ and set $V_{i}\coloneqq V$. Otherwise, if no such $V$ exists that satisfies the aforementioned property, then we just set $V_{i}=T$. Let $T^{\prime}\coloneqq\cap_{i}V_{i}$. By Proposition 5.1, the set ${\rm{bc}}(X_{T^{\prime}}/T^{\prime})$ is finite. For each open subset $T_{0}\subset T^{\prime}$, let $\mathcal{S}_{T_{0}}$ be the set obtained by restricting each element of ${\rm{bc}}(X_{T^{\prime}}/T^{\prime})$ to $T_{0}$. By the finiteness of ${\rm{bc}}(X_{T^{\prime}}/T^{\prime})$ (see Proposition 5.1), there exists a smooth open subset $T_{0}\subset T$ with the following properties:
  1. (1)
    for each $Y/T_{0}\in\mathcal{S}_{T_{0}}$, if $E$ is an exceptional divisor of the birational contraction $X_{T_{0}}\dasharrow Y/T_{0}$, then $E_{\bar{K}}$ is also an exceptional divisor of the birational contraction $X_{\bar{K}}\dasharrow Y_{\bar{K}}$, and
  2. (2)
    if $Y^{\prime}/V\in{\rm{bc}}(X_{V}/V)$ for some $V\subset T$, then there exists $Y/T_{0}\in\mathcal{S}_{T_{0}}$ such that $Y^{\prime}_{\bar{K}}\simeq Y_{\bar{K}}$.
We claim that the desired property holds for ${\rm{bc}}(X_{T_{0}}/T_{0})$. The following modifications are all inside of ${\rm{bc}}(X_{T_{0}}/T_{0})$ and thus we can freely use the minimal model program.
Step 2. Take any $Y^{\prime}/V\in{\rm{bc}}(X_{V}/V)$ for some $V\subset T_{0}$. Then there exists $Y^{(1)}/T_{0}\in\mathcal{S}_{T_{0}}$ such that $Y^{\prime}_{\bar{K}}\simeq Y^{(1)}_{\bar{K}}$ by (2). Let $h_{1}:Y^{(1)}_{V}\dasharrow Y^{\prime}$ be the natural biratoinal map. By $Y^{\prime}_{\bar{K}}\simeq Y^{(1)}_{\bar{K}}$, we see that $h$ is an isomorphism over an open subset $\Omega\subset T_{0}$. By (1), we see that $h_{1}$ is a birational contraction. Indeed, if $E$ is a divisor on $Y^{\prime}$ which is contracted by $h_{1}^{-1}$, then $E$ is also a divisor on $X_{T_{0}}$ which is contracted by $X_{T_{0}}\dasharrow Y^{(1)}$. Hence, (1) implies that $E$ is a horizontal divisor over $T_{0}$. This contradicts with $Y^{\prime}_{\bar{K}}\simeq Y^{(1)}_{\bar{K}}$.
Step 3. Let ${\rm{Exc}}(h_{1})\subset Y_{V}^{(1)}$ be the exceptional divisor. We also denote the corresponding divisor on $Y^{(1)}$ by ${\rm{Exc}}(h_{1})$. Note that ${\rm{Exc}}(h_{1})\cap Y^{(1)}_{\Omega}$$\emptyset$. Let $Y^{(2)}\to Y^{(1)}$ be a small $\mathbb{Q}$-factorial modification. Let $E^{(2)}$ be the strict transform of ${\rm{Exc}}(h_{1})$. Let $Y^{(2)}\dasharrow Y^{(3)}$ be the canonical model of $E^{(2)}$ over $Y^{(1)}$. In particular, if $E^{(3)}$ is the strict transform of $E^{(2)}$, then $E^{(3)}$ is a $\mathbb{Q}$-Cartier divisor. Note that we still have $Y^{(3)}_{\Omega}\simeq Y^{(1)}_{\Omega}\simeq Y^{\prime}_{\Omega}$, and $Y^{(3)}_{V}\dasharrow Y^{\prime}$ is a birational contraction.
[…]
Step 5. Let $H^{\prime}$ be a general very ample/V/V divisor on $Y^{\prime}$. Let $H^{(4)}$ be the strict transform of $H^{\prime}$ on $Y^{(4)}$. Note that $H^{(4)}$ may not be $\mathbb{Q}$-Cartier. Let $Y^{(5)}\to Y^{(4)}$ be a small $\mathbb{Q}$-factorial modification, and let $H^{(5)}$ be the strict transform of $H^{\prime}$ on $Y^{(5)}$. Let $Y^{(5)}\dasharrow Y^{(6)}$ be the canonical model of $H^{(5)}$ over $T_{0}$. Then, $H^{(6)}$, as the strict transform of $H^{\prime}$ on $Y^{(5)}$, is ample over $T_{0}$. Moreover, we have $Y^{(6)}_{\Omega}\simeq Y^{\prime}_{\Omega}$ by construction. As $h_{5}:Y^{(5)}_{V}\dasharrow Y^{\prime}$ is an isomorphism in codimension $1$, and $h_{6}:Y^{(6)}_{V}\dasharrow Y^{\prime}$ is an isomorphism on the locus where $h_{5}$ is an isomorphism, we see that $h_{6}$ is still an isomorphism in codimension $1$.
Note that $Y^{(6)}/T_{0}\in{\rm{bc}}(X_{T_{0}}/T_{0})$. We claim that the natural map $h_{6}$ is an isomorphism, and thus completes the proof. Let $p:W\to Y^{(6)}_{V}$ and $q:W\to Y^{\prime}$ be birational morphisms such that $q\circ p^{-1}=h_{6}$. As $h_{6}$ is an isomorphism in codimension $1$, we have $p^{*}H^{(6)}_{V}=q^{*}H^{\prime}$ by the negativity lemma (see [BCHM10]). Therefore, $h_{6}$ is an isomorphism. This completes the proof. ∎
5.2. MMP in a family
In this subsection, we prove Theorem 1.4 and Theorem 1.6. Recall that
[…]
is a subset of ${\rm{bc}}(X/T)$
Lemma 5.3.
Let $X\to T$ be a projective fibration.
  1. (1)
    Let $X\dashrightarrow Y/T$ be a partial MMP/T/T consisting of divisorial contractions, flipping contractions and flips. Suppose that the natural map $N^{1}(X/T)\to N^{1}(X_{t})$ is surjective for very general $t\in T$. Then there exists an open subset $T_{0}\subset T$ such that the natural map
    $N^{1}(Y/T_{0})\to N^{1}(Y_{t})$
    is surjective for very general $t\in T_{0}$.
  2. (2)
    Assume that $X$ is of Fano type over $T$. Suppose that the natural map $N^{1}(X/T)\to N^{1}(X_{t})$ is surjective for very general $t\in T$. Then there exists an open subset $T_{0}\subset T$ such that for any open subset $V\subset T_{0}$, if $Y/V\in{\rm{bc}}(X_{V}/V)_{\rm MMP}$, then the natural map
    $N^{1}(Y/V)\to N^{1}(Y_{t})$
    is surjective for very general $t\in V$.
Proof.
We divide the proof into several steps.
Step 1. First, we prove statement (1). Shrinking $T$, we can assume that each step of the above MMP/T/T is not a constant map over the generic point of $T$. We claim that the desired property is preserved in each step of this MMP, and thus it holds for $Y/T$.
If $g:X\to X_{1}/T$ is a birational extremal contraction (divisorial or flipping contraction), then we have the commutative diagram
$N^{1}(X/T)$$N^{1}(X_{1}/T)$$N^{1}(X_{t})$$N^{1}(X_{1,t}).$$\scriptstyle{g^{*}}$$\scriptstyle{g_{t}^{*}}$
For a very general $t$, if $B$ is a $\mathbb{Q}$-Cartier divisor on $X_{1,t}$, then $g_{t}^{*}B$ is a $\mathbb{Q}$-Cartier divisor on $X_{t}$ and thus there exists a $\mathbb{Q}$-Cartier divisor $D$ on $X$ such that $D|_{X_{t}}\equiv g_{t}^{*}B$. In particular, we have $D\equiv 0/X_{1}$ as the relative Picard number $\rho(X/X_{1})=1$. This implies that there exists a $\mathbb{Q}$-Cartier divisor $D^{\prime}$ on $X_{1}$ such that $D=g^{*}D^{\prime}$. Thus we have $D^{\prime}|_{X_{1,t}}\equiv B$.
If $g:X\to X_{1}/T$ is a flipping contraction with the flip $g^{+}:X^{+}\to X_{1}/T$, then we claim that for very general $t\in T$, $g_{t}:X_{t}\to X_{1,t}$ is also an $H_{t}$-flipping contraction. Assume that $H$ is a Cartier divisor on $X$ which is anti-ample over $X_{1}$. As $g_{t}$ is a small morphism, it suffices to show $\rho(X_{t}/X_{1,t})=1$. If $B$ is a $\mathbb{Q}$-Cartier divisor on $X_{t}$, then we can assume that there exists a $\mathbb{Q}$-Cartier divisor $D$ on $X$ such that $D_{t}\equiv B$ by the above argument. Hence, we have $D-cH\equiv 0/T$ for some $c\in\mathbb{Q}$. In particular, there exists a $\mathbb{Q}$-Cartier divisor $D^{\prime}$ on $X_{1}$ such that $D-cH=g^{*}D^{\prime}$. Thus we have $D_{t}-cH_{t}=g_{t}^{*}D^{\prime}_{t}$ for the $\mathbb{Q}$-Cartier divisor $D^{\prime}_{t}$ on $X_{1,t}$ which shows the desired claim. As $g_{t}^{+}$ is also a small morphism with $H_{t}^{+}$-ample over $X_{1,t}$, we see that $g_{t}^{+}:X^{+}_{t}\to X_{1,t}$ is the flip of $g_{t}$. If $B^{+}$ is a $\mathbb{Q}$-Cartier divisor on $X_{t}^{+}$, then there exists $r\in\mathbb{Q}$ such that $B^{+}-rH^{+}_{t}\equiv 0$. Hence, there exists a $\mathbb{Q}$-Cartier divisor $B_{1}$ on $X_{1,t}$ such that $(g_{t}^{+})^{*}B_{1}=B^{+}-rH^{+}_{t}$. By the previous argument, there exists a $\mathbb{Q}$-Cartier divisor $D_{1}$ on $X_{1}$ such that $D_{1,t}\equiv B_{1}$. If $D^{+}=(g^{+})^{*}D_{1}$, then we have $D^{+}_{t}+rH^{+}_{t}\equiv B^{+}$. This shows the claim.
Step 2. In this step, we explain that (2) can be reduced to a fixed MMP.
By Proposition 5.2, after shrinking $T$, we can assume that for any open subset $V\subset T$ and $Y/V\in{\rm{bc}}(X_{V}/V)$, there exists some $Y^{\prime}/T\in{\rm{bc}}(X/T)$ such that $Y\simeq Y^{\prime}_{V}/V$. Consider all $Y^{\prime}/T\in{\rm{bc}}(X/T)$ such that there exists an open subset $U\subset T$ with $Y^{\prime}_{U}\in{\rm{bc}}(X_{U}/U)_{\mathrm{MMP}}$. Since ${\rm{bc}}(X/T)$ is a finite set by Proposition 5.1, after shrinking $T$, we may assume that for any open subset $V\subset T$ and any $Y/V\in{\rm{bc}}(X_{V}/V)_{\mathrm{MMP}}$, there exists some $Y^{\prime}/T\in{\rm{bc}}(X/T)_{\mathrm{MMP}}$ such that $Y\simeq Y^{\prime}_{V}/V$.
For $V\subset T_{0}\subset T$, by the natural maps
$N^{1}(Y^{\prime}_{T_{0}}/T_{0})\to N^{1}(Y/V)\to N^{1}(Y_{t}),t\in V,$
it suffices to find an open subset $T_{0}\subset T$ such that for each $Y^{\prime}/T\in{\rm{bc}}(X/T)_{\rm MMP}$, the natural map $N^{1}(Y^{\prime}_{T_{0}}/T_{0})\to N^{1}(Y_{t})$ is surjective for very general $t\in T_{0}$. As ${\rm{bc}}(X/T)_{\rm MMP}$ is a finite set, it suffices to show this property for a fixed $Y^{\prime}/T\in{\rm{bc}}(X/T)_{\rm MMP}$.
Step 3. In this step, we show the statement (2).
If $X\dashrightarrow Y^{\prime}$ only consists of divisorial contractions, flipping contractions and flips, then the claim follows from statement (1).
Suppose that $h:Z\to Y^{\prime}/T$ is a birational contraction between varieties of Fano type/T/T and $N^{1}(Z/T)\to N^{1}(Z_{t})$ is surjective for very general $t\in T$. Note that $h$ corresponds to the contraction of curves belonging to a face $F$ of the Mori cone ${\rm{NE}}(Z/Y^{\prime})$. We can contract an extremal ray of $F$ to obtain a morphism $h^{\prime}:Z\to Z^{\prime}$ and thus $h$ factors through $h^{\prime}$. By repeating this process, we can further assume that $h$ is an extremal contraction. Then the desired property follows from Step 1. ∎
Using Lemma 5.3, we now prove the claim concerning fiber-wise small $\mathbb{Q}$-factorial modifications.
Proof of Theorem 1.4.
Let $h:Y\to X$ be a log resolution of $(X,\Delta)$. Shrinking $T$, we can assume that $Y\to T$ is a smooth morphism. As $X_{s},s\in S$ has rational singularities with $h^{i}(X_{s},\mathcal{O}_{X_{s}})=0,i=1,2$, we can assume that the same property still holds for $Y_{s},s\in S$ after shrinking $T$. By Theorem 1.3 (2), there exits a generically finite base change $T^{\prime}\to T$ and an open subset $T^{\prime}_{0}\subset T^{\prime}$ such that $N^{1}(X_{T_{0}^{\prime}}/T_{0}^{\prime})\to N^{1}(X_{t^{\prime}})$ is surjective for each $t^{\prime}\in T_{0}^{\prime}$. Shrinking $T_{0}^{\prime}$, we can assume that $T_{0}^{\prime}$ is a smooth variety. Thus $Y_{T^{\prime}_{0}}$ is still a smooth variety. Applying Theorem 1.3 (2) to $Y_{T^{\prime}_{0}}\to T^{\prime}_{0}$, there exits a generically finite base change $T^{\prime\prime}\to{T^{\prime}_{0}}$ and an open subset $T_{0}^{\prime\prime}\subset T^{\prime\prime}$ such that $N^{1}(Y_{T_{0}^{\prime\prime}}/T_{0}^{\prime\prime})\to N^{1}(Y_{t^{\prime\prime}})$ is surjective for each $t^{\prime\prime}\in T_{0}^{\prime\prime}$. Note that we still have the surjectivity of the map $N^{1}(X_{T_{0}^{\prime\prime}}/T_{0}^{\prime\prime})\to N^{1}(X_{t^{\prime\prime}}),t^{\prime\prime}\in T_{0}^{\prime\prime}$. Shrinking $T_{0}^{\prime\prime}$, we can assume that there still exists a $\mathbb{Q}$-divisor $\Delta^{\prime\prime}$ such that $(X_{T_{0}^{\prime\prime}},\Delta^{\prime\prime})$ has klt singularities. Thus, we can replace $T^{\prime\prime}_{0}$ by $T$ and assume that
$N^{1}(X/T)\to N^{1}(X_{t})\text{~and~}N^{1}(Y/T)\to N^{1}(Y_{t})$
are surjective for each $t\in T$.
Now we proceed with the standard process of small $\mathbb{Q}$-factorial modifications. Let $K_{Y}+\Delta_{Y}=h^{*}(K_{X}+\Delta)+E$ such that ${\rm{Supp}}E={\rm{Supp}}{\rm{Exc}}(h)$ and $(Y,\Delta_{Y})$ has klt singularities. Shrinking $T$, we can assume that $h$ is a fiber-wise resolution such that the same property of $K_{Y_{t}}+\Delta_{Y_{t}}=h_{t}^{*}(K_{X_{t}}+\Delta_{t})+E_{t}$ holds fiber-wise, where $\Delta_{Y_{t}}=(\Delta_{Y})|_{Y_{t}}$. By [BCHM10], there exists a $(K_{Y}+\Delta_{Y})$-MMP over $X$ which terminates and only $E$ is contracted in this process. Denote this MMP/X/X by
$g:Y=Y_{0}\dashrightarrow Y_{1}\dashrightarrow\cdots\dashrightarrow Y_{n}=W/X.$
Moreover, $g$ only consists of divisorial contractions, flipping contractions and flips. Shrinking $T$, we can assume that if $Y_{j}\to Y_{j+1}$ (resp. $Y_{j+1}\to Y_{j}$) is a divisorial/small contraction (resp. small contraction), then for each $t\in T$, $Y_{j,t}\to Y_{j+1,t}$ (resp. $Y_{j+1,t}\to Y_{j,t}$) is still a divisorial/small contraction (resp. small contraction).
Note that $W\to X$ is a small $\mathbb{Q}$-factorial modification of $X$. We claim that it is also a small $\mathbb{Q}$-factorial modification of each fiber after possible shrinking $T$ again.
By Lemma 5.3 (1), we have a surjection $N^{1}(Y_{j}/T)\to N^{1}(Y_{j,t})$ for $0\leq j\leq n$ over very general point $t\in T$. By Theorem 1.3 (1), after shrinking $T$, we can assume that $N^{1}(Y_{j}/T)\to N^{1}(Y_{j,t})$ is an isomorphism for each $t\in T$ and $0\leq j\leq n$.
Suppose that $Y_{j}\to Y_{j+1}$ is a $(K_{Y_{j}}+\Delta_{Y_{j}})$-negative extremal contraction (divisorial or flipping contraction). By Lemma 4.2, after shrinking $T$, we have the short exact sequences
$\begin{split}0\to&N^{1}(Y_{j+1}/T)\to N^{1}(Y_{j}/T)\to N^{1}(Y_{j}/Y_{j+1})\to 0,~0\leq j\leq n\\ 0\to&N^{1}(Y_{j+1,t})\to N^{1}(Y_{j,t})\to N^{1}(Y_{j,t}/Y_{j+1,t})\to 0,~t\in T.\end{split}$
By the Five Lemma, the commutativity of the following diagram
[…]
implies that $N^{1}(Y_{j}/Y_{j+1})\simeq N^{1}(Y_{j,t}/Y_{j+1,t})$. In particular, $Y_{j,t}\to Y_{j+1,t}$ is still a $(K_{Y_{j,t}}+\Delta_{Y_{j,t}})$-negative extremal contraction. Hence, $Y_{j,t}\to Y_{j+1,t}$ is a divisorial (resp. flipping) contraction if and only if $Y_{j}\to Y_{j+1}$ is a divisorial (resp. flipping) contraction.
Suppose that $Y_{j}\to Y_{j+1}$ is a $(K_{Y_{j}}+\Delta_{Y_{j}})$-negative flipping contraction with flip $Y_{j+2}\to Y_{j+1}$. As $(K_{Y_{j+2,t}}+\Delta_{Y_{j+2,t}})$ is ample over $Y_{j+1}$, we see that $Y_{j+2,t}\to Y_{j+1,t}$ is the flip of the flipping contraction $Y_{j,t}\to Y_{j+1,t}$.
The above shows that for each $t\in T$,
$g_{t}:Y_{t}=Y_{0,t}\dashrightarrow X_{1,t}\dashrightarrow\cdots\dashrightarrow Y_{n,t}=W_{t}/X_{t}$
is also a $(K_{Y_{t}}+\Delta_{Y_{t}})$-MMP over $X_{t}$ which terminates. Thus, the exceptional divisor $E_{t}$ is contracted. As ${\rm{Supp}}E_{t}={\rm{Supp}}{\rm{Exc}}(h_{t})$, we see that $W_{t}\to X_{t}$ is a small $\mathbb{Q}$-factorial modification.
Finally, by Theorem 1.3, after shrinking $T$, for any open subset $U\subset T$, the natural restriction maps
$N^{1}(W/T)\to N^{1}(W_{U}/U)\to N^{1}(W_{t}),\quad N^{1}(X/T)\to N^{1}(X_{U}/U)\to N^{1}(X_{t})$
are isomorphisms for any $t\in U$. ∎
The following proposition shows that if $X$ is $\mathbb{Q}$-factorial and of Fano type over $T$, then any of its birational contraction models can be obtained via a partial MMP over $T$.
Proposition 5.4.
If $X$ is a $\mathbb{Q}$-factorial variety of Fano type over $T$, then
${\rm{bc}}(X/T)={\rm{bc}}(X/T)_{\rm MMP}.$
Proof.
By definition, we have ${\rm{bc}}(X/T)\supset{\rm{bc}}(X/T)_{\rm MMP}$.
For the inverse inclusion, let $Y/T\in{\rm{bc}}(X/T)$. We show that $Y$ can be obtained from $X$ via an MMP/T/T followed by a birational contraction. To be precise, we need to show that the birational contraction $X\dashrightarrow Y/T$ corresponding to $Y$ can be factored into an MMP/T/T followed by a birational contraction up to an isomorphism of $Y/T$.
Let $A^{\prime}$ be an ample/T/T divisor on $Y$, and let $A$ be its strict transform on $X$. Let $p:W\to X$ and $q:W\to Y$ be birational morphisms such that $q\circ p^{-1}$ is the map $X\dashrightarrow Y$. As $X$ is a $\mathbb{Q}$-factorial variety, we can write
$p^{*}A+E=q^{*}A^{\prime}+F,$
where $E,F$ are effective $q$-exceptional divisors. Let $B=p_{*}(p^{*}A+E)$ be a $\mathbb{Q}$-Cartier divisor on $X$, then
$p^{*}B+E_{1}=p^{*}A+E+E_{2},$
where $E_{1},E_{2}$ are effective $p$-exceptional divisors. Since $X\dashrightarrow Y$ is a birational contraction, $E_{1},E_{2}$ are also $q$-exceptional divisors. We have
$p^{*}B+E_{1}=p^{*}A+E+E_{2}=q^{*}A^{\prime}+F+E_{2}.$
Since $A^{\prime}$ is ample/T/T and both $F$ and $E_{2}$ are $q$-exceptional divisors, the divisors $q^{*}A^{\prime}$ and $F+E_{2}$ form the positive and negative parts, respectively, of the Nakayama-Zariski decomposition of $p^{*}B+E_{1}$ over $T$. Let $X\dashrightarrow Z/T$ be a $B$-MMP/T/T such that $B_{Z}$, the strict transform of $B$, is semi-ample/T/T. Replacing $W$ by a higher model, we can assume that there exists a birational morphism $r:W\to Z$. We have
$p^{*}B=r^{*}B_{Z}+E_{Z},$
where $E_{Z}$ is an effective $r$-exceptional divisor. Since $E_{1}$ is also $r$-exceptional, $p^{*}B+E_{1}$ admits the Nakayama-Zariski decomposition whose positive part is $r^{*}B_{Z}$ and the negative part is $E_{Z}+E_{1}$. By the uniqueness of the Nakayama-Zariski decomposition, we have $E_{Z}+E_{1}=F+E_{2}$ and $r^{*}B_{Z}=q^{*}A^{\prime}$. As $A^{\prime}$ is ample/T/T, the natural birational map $g:Z\dashrightarrow Y/T$ is indeed a morphism. ∎
Lemma 5.5.
Let $X$ be of Fano type over $T$. Then, up to a generically finite base change of $T$, there exists an open subset $T_{0}\subset T$ such that the following properties hold:
  1. (1)
    for any open subset $V\subset T_{0}$, if $Y/V\in{\rm{bc}}(X_{V}/V)$, then the natural maps
    $N^{1}(Y/V)\to N^{1}(Y_{t}),\quad{\rm{Nef}}(Y/V)\to{\rm{Nef}}(Y_{t}),\quad{\rm{NE}}(Y_{t})\to{\rm{NE}}(Y/V)$
    are isomorphisms for any $t\in V$;
  2. (2)
    for any $T_{0}\in{\rm{bc}}(X/T_{0}), YY$ is flat over $T_{0}$, and $Y_{t}$ is an irreducible normal variety for each $t\in T_{0}$.
  3. (3)
    if $V\subset T_{0}$ is an open subset and $g:Y\to Z/V$ is a contraction (may not be birational) for some $Y\in{\rm{bc}}(X_{V}/V)$, then $g_{t}:Y_{t}\to Z_{t}$ is still a contraction for each $t\in V$. Moreover, if $g$ is a divisorial contraction (resp. a small contraction, an extremal contraction, a Mori fiber space) if and only if $g_{t},t\in V$ is a divisorial contraction (resp. a small contraction, an extremal contraction, a Mori fiber space).
Proof.
By Theorem 1.4, there exists a generically finite base change $T^{\prime}\to T$ and an open subset $T_{0}^{\prime}\subset T^{\prime}$ such that $X_{T_{0}^{\prime}}$ admits a small $\mathbb{Q}$-factorial modification $W\to X_{T_{0}^{\prime}}$ which is also a fiber-wise small $\mathbb{Q}$-factorial modification over $T_{0}^{\prime}$. Moreover, the natural map $N^{1}(W/U)\to N^{1}(W_{t}),t\in U$ is an isomorphism, where $U\subset T_{0}^{\prime}$ is an open subset. Note that $W$ is of Fano type over $T_{0}^{\prime}$ after shrinking $T_{0}^{\prime}$. By Proposition 5.4, we have
${\rm{bc}}(W/T_{0}^{\prime})_{\rm MMP}={\rm{bc}}(W/T^{\prime}_{0})={\rm{bc}}(X_{T_{0}^{\prime}}/T_{0}^{\prime}).$
Replacing $W$ by $X$ and $T_{0}^{\prime}$ by $T$, we may assume that $Y/V\in{\rm{bc}}(X_{V}/V)_{\rm{MMP}}$ and that the natural map $N^{1}(X_{V}/V)\to N^{1}(X_{t})$ is an isomorphism for any $t\in V\subset T$. Shrinking $T$, we can assume that Proposition 5.2 also holds. By Proposition 5.3 (2), after shrinking $T$, we can assume that for any open subset $V\subset T$ and $Y/V\in{\rm{bc}}(X_{V}/V)_{\rm MMP}$, $N^{1}(Y/V)\to N^{1}(Y_{t})$ is surjective for very general $t\in V$. By Theorem 1.3 (1) and the finiteness of ${\rm{bc}}(X_{V}/V)_{\rm MMP}$, after shrinking $T$, we can assume that
$N^{1}(Y/V)\to N^{1}(Y_{t}),t\in V$
is an isomorphism for any open subset $V\subset T$ and any $Y/V\in{\rm{bc}}(X_{V}/V)_{\rm MMP}$. By Corollary 4.6, after shrinking $T$, we can assume that for any $Y/T\in{\rm{bc}}(X/T)_{\rm MMP}$ and any open subset $V\subset T$, the natural maps
${\rm{Nef}}(Y_{V}/V)\to{\rm{Nef}}(Y_{t}),\quad{\rm{NE}}(Y_{t})\to{\rm{NE}}(Y/V)$
are isomorphisms for each $t\in V$. This shows (1). Since each step in the above process remains valid after shrinking $T$, the statement (1) continues to hold after shrinking $T$.
For (2), after shrinking $T$, we can assume that for any $T\in{\rm{bc}}(X/T), YY$ is flat over $T$ by generic flatness, and $Y_{t},t\in T$ is an irreducible normal variety as $\{p\in T\mid Y_{p}\text{~is normal}\}$ is a constructible set. Note that, in the above set, $p$ is not restricted to closed points, and the generic point of $T$ belongs to this set as $Y$ is normal.
To show (3), first note that by (1), we have ${\rm{Nef}}(Y/T)\simeq{\rm{Nef}}(Y_{V}/V)$ for any $Y/T\in{\rm{bc}}(X/T)$ and open subset $V\subset T$. As ${\rm{Nef}}(Y/T)$ is a polyhedral cone, there are only finitely many contractions (not just birational contractions). Moreover, if $Y^{\prime}\to Z^{\prime}/V$ is a contraction for some $Y^{\prime}/V\in{\rm{bc}}(X_{V}/V)$, then there exist $Y/T\in{\rm{bc}}(X/T)$ and a contraction $g:Y\to Z/T$ such that $g_{V}=h$. As ${\rm{NE}}(Y/T)\simeq{\rm{NE}}(Y_{V}/V)\simeq{\rm{NE}}(Y_{t})$, we see that $g,g_{V},g_{t}$ are extremal contractions if one of them is an extremal contraction. Shrinking $T$ further, we can assume that $g$ is a divisorial contraction (resp. a small contraction, a Mori fiber space) if and only if $g_{t},t\in V$ is a divisorial contraction (resp. a small contraction, a Mori fiber space). ∎
Proof of Theorem 1.6.
By Theorem 1.2, after shrinking $T$, we can assume that $X/T$ is of Fano type. After replacing $T$ by a generically finite base change and possibly shrinking it, we may assume that the properties listed in Lemma 5.5 hold.
First, by Proposition 5.1 and Proposition 5.2, it suffices to consider a fixed birational contraction $g:X^{\prime}\dashrightarrow Y/U$. Take birational morphisms $p\colon W\to X^{\prime}$ and $q\colon W\to Y$ such that $g=q\circ p^{-1}$. Shrinking $U$, we may assume that for any $t\in U$, the induced map
$g_{t}\colon X^{\prime}_{t}\dashrightarrow Y_{t}$
is still a birational contraction, and that $p_{t}$ and $q_{t}$ are birational morphisms. Moreover, we may further assume that if a divisor $E$ on $W$ is $p$-exceptional (resp. $q$-exceptional), then for every $t\in U$, the divisor $E|_{W_{t}}$ is $p_{t}$-exceptional (resp. $q_{t}$-exceptional). By construction, for any $\mathbb{R}$-Cartier divisor $\mathcal{D}$ on $X^{\prime}$, we have
$p^{*}\mathcal{D}=q^{*}(g_{*}\mathcal{D})+E,$
where $E$ is $q$-exceptional. Hence, we have
$p_{t}^{*}\mathcal{D}|_{X_{t}^{\prime}}\sim_{\mathbb{R}}q_{t}^{*}((g_{*}\mathcal{D})|_{Y_{t}})+E|_{W_{t}}.$
As $g_{t}$ is a birational contraction and $E|_{W_{t}}$ is $q_{t}$-exceptional, we see
$g_{t,*}(\mathcal{D}|_{X_{t}^{\prime}})\sim_{\mathbb{R}}(g_{*}\mathcal{D})|_{Y_{t}}.$
Next, we show that an MMP of the total space is an MMP of each fiber of the same type. For any $\mathcal{D}\in N^{1}(X/T)$, let
$X=X_{0}\dasharrow X_{1}\dasharrow\cdots\dashrightarrow X_{n-1}\dasharrow X_{n}\to X_{n+1}$
be a $\mathcal{D}$-MMP over $T$, where $X_{i}\dasharrow X_{i+1},i=0,\ldots,n-1$ are birational contractions, and $X_{n}\to X_{n+1}$ is either a contraction induced by the semi-ample/T/T divisor $\mathcal{D}_{X_{n}}$ or a Mori fiber space when $\mathcal{D}$ is not pseudo-effective/T/T. Thus, we have $X_{i}/T\in{\rm{bc}}(X/T)_{\rm MMP},0\leq i\leq n$, and $X_{n+1}/T\in{\rm{bc}}(X/T)_{\rm MMP}$ if $X\dasharrow X_{n+1}$ is birational. To be precise, this means that the above natural birational contraction $X\dasharrow X_{i}/T$ belongs to ${\rm{bc}}(X/T)_{\rm MMP}$. By Lemma 5.5 (3), if $g:X_{i}\dasharrow X_{i+1}$ is a divisorial contraction (resp. a small contraction, an extremal contraction, a Mori fiber space) if and only if $g_{t}$ is a divisorial contraction (resp. a small contraction, an extremal contraction, a Mori fiber space) for each $t\in T$. Hence, for each $t\in T$,
$X_{t}=X_{0,t}\dasharrow X_{1,t}\dasharrow\cdots\dasharrow X_{n,t}\to X_{n+1,t}$
is a $\mathcal{D}_{t}$-MMP on $X_{t}$ of the same type.
Conversely, we show that an MMP of the fiber is induced from an MMP/T/T of the total space. Suppose that
$X_{t}=Y_{0}\dasharrow Y_{1}\dasharrow\cdots\dasharrow Y_{n}\to W$ (5.2)
is a $D$-MMP on $X_{t}$. By Lemma 5.5 (1), there exists $\mathcal{D}\in N^{1}(X/T)$ such that $\mathcal{D}_{t}\sim_{\mathbb{R}}D$. If $\sigma_{0}:Y_{0}\to Y_{1}$ is a divisorial contraction (resp. a small contraction, an extremal contraction, a Mori fiber space), then by Lemma 5.5 (1), there exists a contraction $g_{0}:X_{0}\to X_{1}$ that contracts the same extremal ray as $\sigma_{0}$, and thus $g_{0,t}=\sigma_{0}$. By Lemma 5.5 (3), $g_{0}$ is still a divisorial contraction (resp. a small contraction, an extremal contraction, a Mori fiber space). When $\sigma_{0}$ is a flipping contraction, let $\sigma_{1}:Y_{2}\to Y_{1}$ be its flip. Let $g_{1}:X_{2}\to X_{1}$ be the flip of the flipping contraction $g_{0}:X_{0}\to X_{1}$. By Lemma 5.5 (2) (3), $g_{1,t}:X_{2,t}\to X_{1,t}$ is an extremal contraction between normal varieties. Moreover, $\mathcal{D}_{X_{2}}|_{X_{2,t}}$ is ample over $X_{1,t}$ as $\mathcal{D}_{X_{2}}$ is ample over $X_{1}$, where $\mathcal{D}_{X_{2}}$ is the strict transform of $\mathcal{D}$ on $X_{2}$. Therefore, $g_{1,t}$ is exactly $\sigma_{1}$. Repeating this process, we obtain a $\mathcal{D}$-MMP on $X/T$ whose restriction to the fiber $X_{t}$ is exactly (5.2). Moreover, this $\mathcal{D}$-MMP terminates when (5.2) terminates by Lemma 5.5 (1) (3).
Finally, if $Y/U\in{\rm{bc}}(X_{U}/U)_{\mathrm{MMP}}$, then (3) directly follows from Lemma 5.5 (1). In general, by Theorem 1.4, after a generic base change of $T$ and shrinking it, there exists a fiber-wise small $\mathbb{Q}$-factorial modification $W\to X/T$. By Proposition 5.4, we have $Y/U\in{\rm{bc}}(W_{U}/U)_{\rm MMP}$. Then by the same argument as before, (3) follows from Lemma 5.5 (1). ∎
5.3. Deformation of effective cones
As an application of Theorem 1.6, we show the following result on the deformation invariance of effective cones.
Proposition 5.6.
Let $f:X\to T$ be a projective fibration. Suppose that there exists a Zariski dense subset $S\subset T$ such that the fibers $X_{s}$ for $s\in S$ satisfy one of the cases in Theorem 1.2. Then, up to a generically finite base change of $T$, there exists a non-empty open subset $T_{0}\subset T$, such that for any $U\subset T_{0}$, we have
${\rm{Eff}}(X_{T_{0}}/T_{0})\simeq{\rm{Eff}}(X_{U}/U)\simeq{\rm{Eff}}(X_{t}),t\in U$
under the natural restriction maps.
Proof.
After replacing $T$ by a generically finite base change and possibly shrinking it, we may assume that Theorem 1.6 holds. By Theorem 1.2, after shrinking $T$, we can assume that $X/T$ is of Fano type. If $\mathcal{D}\in{\rm{Eff}}(X/T)$, then we can run a $\mathcal{D}$-MMP/T/T, $X\dashrightarrow X_{n}$, which terminates to $X_{n}$ such that $\mathcal{D}_{X_{n}}$ is nef /T/T. By Theorem 1.6 (2), it is also a $\mathcal{D}_{t}$-MMP on $X_{t}$ for any $t\in T$. This shows that under the identification $N^{1}(X/T)\simeq N^{1}(X_{t})$ (see Theorem 1.6 (4)), we have
${\rm{Eff}}(X/T)\subset{\rm{Eff}}(X_{t}).$
Conversely, take $D\in{\rm{Eff}}(X_{t})$. Let $X_{t}\dasharrow Y$ be a $D$-MMP such that $D_{Y}$ is nef on $Y$. Then by Theorem 1.6 (3), it is the restriction of a $\mathcal{D}$-MMP/T/T on $X$, where $\mathcal{D}$ is a divisor on $X$ such that $\mathcal{D}_{t}]=[D]$. Hence we have $\mathcal{D}]\in{\rm{Eff}}(X/T)$. This show the inverse inclusion ${\rm{Eff}}(X/T)\supset{\rm{Eff}}(X_{t})$. ∎
5.4. Deformation of movable cones and Mori chamber decompositions
The deformation invariance of movable cones and Mori chamber decompositions is also a consequence of Theorem 1.6.
If $X$ is of Fano type over $T$ and $D]\in{\rm{Mov}}(X/T)=\overline{\rm{Mov}}(X/T)$ is a movable divisor, then there exists a $D$-MMP/T/T, $\phi:X\dashrightarrow Y/T$, which terminates at $Y$. Note that $X\dashrightarrow Y$ is isomorphic in codimension $1$ and if $D_{Y}$ is the strict transform of $D$ on $Y$, then $D_{Y}]\in{\rm{Nef}}(Y/T)$. Conversely, if $\phi:X\dashrightarrow Y$ is isomorphic in codimension $1$, then $\phi^{-1}_{*}{\rm{Nef}}(Y/T)\cap N^{1}(X/T)\subset{\rm{Mov}}(X/T)$. Here the chamber $\phi^{-1}_{*}{\rm{Nef}}(Y/T)\cap N^{1}(X/T)$ consists of $\mathbb{R}$-Cartier divisors which are strict transforms of nef divisors on $Y/T$.
Therefore, ${\rm{Mov}}(Y)$ admits the following Mori chamber decomposition
${\rm{Mov}}(X/T)=\bigcup_{\begin{subarray}{c}\phi:X\dashrightarrow Y\text{~isom.~in codim.~}1,\\ Y/T\in{\rm{bc}}(X/T)_{\rm MMP}\end{subarray}}\left(\phi^{-1}_{*}{\rm{Nef}}(Y/T)\cap N^{1}(X/T)\right).$ (5.3)
We emphasize that $Y/T\in{\rm{bc}}(X/T)_{\mathrm{MMP}}$ indeed represents the map $\phi$.
This Mori chamber decomposition may differ from
${\rm{Mov}}(X/T)=\bigcup_{\phi:X\dashrightarrow Y\text{~isom.~in codim.~}1}\left(\phi^{-1}_{*}{\rm{Nef}}(Y/T)\cap N^{1}(X/T)\right),$ (5.4)
where all birational maps that are isomorphic in codimension $1$ are considered. Indeed, if $\psi:\tilde{X}\to X$ is a non-isomorphic small $\mathbb{Q}$-factorial modification, then the chamber $\left(\psi_{*}{\rm{Nef}}(\tilde{X}/T)\cap N^{1}(X/T)\right)$ appears in (5.4). However, it does not appear in (5.3) because $X\dashrightarrow\tilde{X}$ does not belong to ${\rm{bc}}(X/T)_{\rm MMP}$. On the other hand, if $X$ is $\mathbb{Q}$-factorial, then Proposition 5.4 implies that the two chamber decompositions are the same. It is the Mori chamber decomposition given in (5.3) that will be considered below. (A similar statement can be obtained for the decomposition in (5.4) by passing to a $\mathbb{Q}$-factorial modification, but the formulation is more involved and will be omitted here.)
Proposition 5.7.
Let $f:X\to T$ be a projective fibration. Suppose that there exists a subset $S\subset T$ such that the fibers $X_{s}$ for $s\in S$ satisfy one of the cases in Theorem 1.2. Then, up to a generically finite base change of $T$, there exists a Zariski open subset $T_{0}\subset T$, such that for any $U\subset T_{0}$, we can identify the Mori chamber decompositions of ${\rm{Mov}}(X_{T_{0}}/T_{0}),{\rm{Mov}}(X_{U}/U)$ and ${\rm{Mov}}(X_{t}),t\in U$ under the natural restriction maps. In particular, we have isomorphisms among movable cones
${\rm{Mov}}(X_{T_{0}}/T_{0})\to{\rm{Mov}}(X_{U}/U)\to{\rm{Mov}}(X_{t}),t\in U$
under the natural restriction maps.
Proof.
After replacing $T$ by a generically finite base change and possibly shrinking it, we may assume that Theorem 1.6 holds. By Theorem 1.2, after shrinking $T$, we can assume that $X/T$ is of Fano type. By Theorem 1.6 (2) and (3), an MMP/T/T on $X$ restricts to an MMP of the same type on each fiber $X_{t}$, and the converse also holds. Therefore, ${\rm{Mov}}(X/T)$ and ${\rm{Mov}}(X_{t})$ shares the same chambers under the identification $N^{1}(X/T)\simeq N^{1}(X_{t})$ (see Theorem 1.6 (4)). ∎
6. Boundedness of birational models of Fano type varieties
Based on the results developed in the previous sections, we present an application to the moduli problem of Fano type varieties.
Proof of Theorem 1.7.
By definition, $\mathcal{S}$ consists of closed fibers of a family of Fano type varieties $X/T$. By Noetherian induction and Theorem 1.4, after replacing $T$ by a generically finite base change and possibly shrinking it, we can assume that there exists a fiber-wise small $\mathbb{Q}$-factorial modification $Y\to X$. Note that ${\rm bcm}(Y_{t})={\rm bcm}(X_{t})$. Replacing $Y$ by $X$, we can assume that $X_{t},t\in T$ are $\mathbb{Q}$-factorial.
After replacing $T$ by a generically finite base change and possibly shrinking it further, we may assume that Theorem 1.6 holds. Note that this replacement does not affect the $\mathbb{Q}$-factoriality of the fibers. By Theorem 1.2 and Noetherian induction, after shrinking $T$, we can assume that $X$ is of Fano type over $T$.
Fix a $X_{t}$ for some $t\in T$, and suppose that $X_{t}\dasharrow Z$ is a birational contraction. As $X_{t}$ is a $\mathbb{Q}$-factorial variety, we have $Z\in{\rm{bc}}(X_{t})_{\rm MMP}$ by Proposition 5.4. By Theorem 1.6 (3), there exists an MMP/T/T, denoted by $X\dasharrow\mathcal{Z}/T$, such that $Z\simeq\mathcal{Z}_{t}$. Note that $\mathcal{Z}/T\in{\rm{bc}}(X/T)$. By Proposition 5.1, ${\rm{bc}}(X/T)$ is a finite set. Thus, ${\rm bcm}(\mathcal{S})$ is bounded by definition. ∎
Remark 6.1.
In contrast to Theorem 1.7, boundedness fails even when considering crepant models of a fixed log pair.
For example, let $X=\mathbb{P}^{2}$, and let $D$ and $B$ be two simple normal crossing curves on $X$ with $p\in D\cap B$. Set $c_{n}=1-\frac{1}{n}$ for $n\in\mathbb{Z}_{>1}$. Suppose that $\pi_{1}:X_{1}={\rm Bl}_{p}X\to X$ is the blowing up of $X$ at $p$ and let $D_{1},B_{1}$ be the strict transforms of $D,B$ on $X_{1}$, respectively. Let $E_{1}$ be the exceptional divisor of $\pi_{1}$. Then we have
$K_{X_{1}}+c_{n}D_{1}+c_{n}B_{1}+(1-\frac{2}{n})E_{1}=\pi_{1}^{*}(K_{X}+c_{n}D+c_{n}B).$
Next, blowing up $D_{1}\cap E_{1}$, we get a variety $X_{2}$ with $D_{2},B_{2}$ the strict transforms of $D_{1},B_{1}$ on $X_{2}$, respectively. We still use $E_{1}$ to denote the strict transform of $E_{1}$ on $X_{2}$, and $E_{2}$ to denote the exceptional divisor of $X_{2}\to X_{1}$. Then we have
$K_{X_{2}}+c_{n}D_{2}+c_{n}B_{2}+(1-\frac{2}{n})E_{1}+(1-\frac{3}{n})E_{2}=\pi_{2}^{*}(K_{X}+c_{n}D+c_{n}B),$
where $\pi_{2}:X_{2}\to X$ is the corresponding morphism. We continue this process by successively blowing up each intersection $D_{i}\cap E_{i}$ at every step. This yields the equation
$K_{X_{m}}+c_{n}D_{m}+c_{n}B_{m}+(1-\frac{2}{n})E_{1}+\cdots+(1-\frac{m+1}{n})E_{m}=\pi_{m}^{*}(K_{X}+c_{n}D+c_{n}B),$
where the notation follows the same meaning as above. In particular, there exists a divisor $E_{n-1}$ whose discrepancy with respect to $(X,c_{n}D+c_{n}B)$ is 0. By [BCHM10], we can extract $E_{n-1}$ alone to get a variety $\theta_{n}:Y_{n}\to X$. Therefore, we have
$K_{Y_{n}}+c_{n}D_{Y_{n}}+c_{n}B_{Y_{n}}=\theta_{n}^{*}(K_{X}+c_{n}D+c_{n}B)$
with $E_{n-1}$ the only exceptional divisor of $\theta_{n}$, where $D_{Y_{n}},B_{Y_{n}}$ are strict transforms of $D,B$ on $Y_{n}$, respectively. In other words, $(Y_{n},c_{n}D_{Y_{n}}+c_{n}B_{Y_{n}})$ is a crepant model of $(X,c_{n}D+c_{n}B)$. Note that $\{E_{n}\mid n\in\mathbb{Z}_{\geq 1}\}$ consists of distinct divisors.
We claim that $\{Y_{n}\mid n\in\mathbb{Z}_{\geq 2}\}$ does not belong to a bounded family. If $\{Y_{n}\mid n\in\mathbb{Z}_{\geq 2}\}$ belongs to a bounded family $\mathcal{Y}\to T$, then without loss of generality, we can assume that there exists a Zariski dense subset $S\subset T$ parametrizing a subset of $\{Y_{n}\mid n\in\mathbb{Z}_{\geq 2}\}$. After shrinking $T$, we can assume that it satisfies the assumptions in Theorem 1.6. Possibly shrinking $T$ further, by Theorem 1.5, we can assume that there exists an effective divisor $\mathcal{E}\subset\mathcal{Y}$ such that $\mathcal{E}_{s_{0}}$ is $\mathbb{Q}$-linearly equivalent to the exceptional divisor of $\mathcal{Y}_{s_{0}}\simeq Y_{i}\to X=\mathbb{P}^{2}$ for a fixed $s_{0}\in S\cap T$. By Theorem 1.6, we can contract $\mathcal{E}$, and obtain the morphism $\Theta:\mathcal{Y}\to\mathcal{X}/T$, which is exactly $\mathcal{Y}_{s_{0}}\simeq Y_{i}\to X=\mathbb{P}^{2}$ over $s_{0}$. By the rigidity of $\mathbb{P}^{2}$, we have $\mathcal{X}_{T^{\prime}}\simeq\mathbb{P}^{2}\times T^{\prime}$, where $T^{\prime}\to T$ is a finite base change ([Har10, Example 5.3.1, Exercise 24.7(c)]). Replacing $T^{\prime}$ by $T$, we can assume that $\mathcal{X}\simeq\mathbb{P}^{2}\times T$. Note that for any $Y_{j}$, there exists a unique morphism $\theta_{j}:Y_{j}\to\mathbb{P}^{2}$, which is the one obtained in the above construction. Indeed, the exceptional curve of $\theta_{j}$ must be an exceptional curve of any other $Y_{j}\to\mathbb{P}^{2}$ by the negativity of the intersection number. From this, we know that if $s_{j}\in S\cap T$ with $\mathcal{Y}_{s_{j}}\simeq Y_{j}$, then $\Theta|_{\mathcal{Y}_{s_{j}}}=\theta_{j}$. Now, by the construction of the divisors $E_{n}$, there exists a rational function $f\in K(X)$ such that $\nu_{E_{i}}(f)=0$ but $\nu_{E_{j}}(f)>0$ for each $j>i$, where $\nu_{E_{n}}$ denotes the valuation corresponding to $E_{n}$. This is because $E_{j},j>i$, is obtained as further blowing-ups over a point of $E_{i}$. Now we take $f$ as a rational function on $\mathcal{X}=\mathbb{P}^{2}\times T$. By the above discussion, the zero set $\mathcal{Z}$ of $f$ on $\mathcal{Y}$ contains $E_{j}=\mathcal{E}|_{s_{j}}$, where $j>i$ (as $\nu_{E_{j}}(f)>0$). By the density of $S$, we have $\mathcal{Z}\supset\mathcal{E}$. However, this contradicts to the fact that $\nu_{E_{i}}(f)=0$ (i.e., $f$ does not vanish on $E_{i}=\mathcal{E}|_{s}$). This shows that the set $\{Y_{n}\mid n\in\mathbb{Z}_{\geq 2}\}$ is not bounded.